### v3

# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

Abstract
Magnetar is a neutron star with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G) which usually manifests as soft gamma-ray repeater (SGR) or anomalous X-ray pulsar (AXP). SGR J1935+2154 is not only one of the most active magnetar detected so far, but also the unique confirmed source of fast radio burst (FRB). Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM) are dedicated to monitor gamma-ray transients all over the sky, including SGR bursts. Here we report the GECAM observation of the burst activity of SGR J1935+2154 from January 2021 to December 2022, which results in a unique and valuable data set for this important magnetar. With a targeted search of GECAM data, 164 bursts from SGR J1935+2154 are detected by GECAM-B while 97 bursts by GECAM-C, including the X-ray burst associated with a fast radio burst (FRB 20221014). We find that both the burst duration and the waiting time between two successive bursts follow lognormal distributions. The period of burst activity is $134\pm 20$ days, thus the burst activity could be generally divided into 4 active episodes over these two years. Interestingly, the hardness ratio of X-ray bursts tends to be softer and more concentrated over these two years, especially during the active episode with FRBs detected.
magnetars - soft gamma-ray repeaters: general - methods: data analysis - techniques
1 Introduction
The soft gamma-ray repeaters (SGRs) and anomalous X-ray pulsars (AXPs) are perceived to originate in the highly magnetized neutron star, namely magnetar (Duncan & Thompson, 1992; van Kerkwijk et al., 1995; Kouveliotou et al., 1998; Banas et al., 1997; Kaspi & Beloborodov, 2017). Magnetars are characterized by slow rotation period ($P\sim 2-12\ \mathrm{s}$), rapidly spin down ($\dot{P}\sim 10^{-13}-10^{-11}\ \mathrm{s\cdot s^{-1}}$) and relatively young age (typically about 1000 yr). So far, 30 magnetars had been detected and 24 of them had been confirmed (Olausen & Kaspi, 2014). Most magnetars can increase persistent radiation and emit bursts/flares simultaneously in the X-/gamma-ray band during an outburst.
Based on their luminosity and duration, the SGR bursts can be divided into three classes (Woods & Thompson, 2006): Short-duration burst, which consists of single or multiple pulses, is the most typical magnetar burst with a burst duration $\sim$0.01 $\mathrm{s}$ and the luminosity around $10^{38}\sim 10^{40}$ $\mathrm{erg\cdot s^{-1}}$; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ($>1$ $\mathrm{s}$) and peak luminosity around $10^{41}\sim 10^{43}$ $\mathrm{erg\cdot s^{-1}}$; Giant flare, the rarest and the most powerful energetic burst, is characterized by a significantly higher luminosity than a typical magnetar burst and a special pulse profile with a hard initial spike and rapidly decaying tail.
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos et al., 2014), while it is worth to note that Xie et al. (2022) found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan et al., 2009) before the this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, $P\sim 3.24\ \mathrm{s}$ and $\dot{P}\sim 1.43\times 10^{-11}\ \mathrm{s\cdot s^{-1}}$, respectively. This indicates a dipole-magnetic field of $B\sim 2.2\times 10^{14}\ \mathrm{G}$ (Israel et al., 2016). It has experienced multiple active episodes from 2014 to 2022 (Younes et al., 2017; Lin et al., 2020b, a; Cai et al., 2022a, b; Xie et al., 2022; Rehan & Ibrahim, 2023, 2024). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which a burst forest was observed (Kaneko et al., 2021). Among these bursts there is an X-ray burst (XRB) (Li et al., 2021; Mereghetti et al., 2020; Ridnaia et al., 2021; Tavani et al., 2020) associated with a fast radio burst, FRB 200428 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020a). After the FRB 20200428 event, GECAM (Li et al., 2021; Xiao et al., 2022a; Zhang et al., 2023; Wang et al., 2024) detected an X-ray burst from SGR J1935+2154 which associated with a radio burst detected by CHIME on 2022 October 14 (Wang et al., 2022a). Later, the Insight–Hard X-ray Modulation Telescope (HXMT) again detected an X-ray burst associated with a radio burst found by the Yunnan 40 m telescope on 2022 October 21 (Li et al., 2022).
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observation data from January 2021 to December 2022. As Paper I of this series work for GECAM observation of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (In prep.), we will give the detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section 2, we present the observation, search process, and report the burst catalog. In Section 3, we present the catalog analysis method and results. Finally, the summary and discussions are given in Section 4.
Please note that the error reported in this work is for 1 sigma, if not stated otherwise.
2 Observation and Burst Search
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, Xiao et al., 2022b), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang et al., 2023) and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang et al., 2024), dedicated to monitor all-sky gamma-ray transients. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 Charged particle detectors (CPDs) while GECAM-C has 12 GRDs and 2 CPDs. All GRDs of GECAM-B and most of GRDs of GECAM-C operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time.
2.1 Search and Identification
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan & Ibrahim, 2023, 2024; Wood & Fermi-GBM Team, 2022). A search pipeline developed by Cai et al. (2021) for gamma-ray bursts (GRBs) is used to do a blind search on GECAM or HXMT daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B, as reported in Xie et al. (2022) and 75 X-ray bursts were found by HXMT (Cai et al., 2022a).
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold, we conduct a targeted sub-threshold search on GECAM daily event-by-event (EVT) data [^1] https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai et al., submitted. Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin et al., 2020b, a; Rehan & Ibrahim, 2023, 2024), we utilize the following spectral templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameters setting is listed in Table 1 which is similar with the Table 1 in Xie et al. (2022). The details of each template are presented below.
The Band function:
$A(E)\sim\begin{cases}E^{\alpha}\exp\left[-\frac{(2+\alpha)E}{E_{\mathrm{p}}}\right],&\text{if }E<\frac{(\alpha-\beta)E_{\mathrm{p}}}{2+\alpha},\\ \\ E^{\beta}\exp(\beta-\alpha)\left[\frac{(\alpha-\beta)E_{\mathrm{p}}}{(2+\alpha)}\right]^{\alpha-\beta},&\text{otherwise,}\end{cases}$ (1)
where $E_{\mathrm{p}}$, in units of keV, is the peak energy of the $\nu F_{\mathrm{\nu}}$ spectrum, $\alpha$/$\beta$ are the low-/high-energy photon indices.
The CPL function:
$A(E)\sim E^{-\alpha}\exp\left[-\frac{(2-\alpha)E}{E_{\mathrm{p}}}\right],$ (2)
where $E_{\mathrm{p}}$, in units of keV, is the peak energy of the $\nu F_{\mathrm{\nu}}$ spectrum, $\alpha$ are the photon index.
The OTTB function:
$A(E)\sim E^{-1}\exp\left(-\frac{E}{kT}\right),$ (3)
where $kT$ is the electron energy, in units of keV.
The Blackbody function:
$A(E)\sim\frac{E^{2}}{\exp(E/kT)-1},$ (4)
where $kT$ is the temperature, in units of keV.
The Powerlaw function:
$A(E)\sim E^{-\alpha},$ (5)
where $\alpha$ are the photon index.
Thousands of trigger (burst candidates) have been found through the targeted search process. We execute a refined analysis for each burst candidate by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools [^2] https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from candidates should adhere to the following criteria (same as Xie et al., 2022):
  1. 1.
    Using the GECAM position history file (posatt) to retrieve the position of satellite and the pointing of detector. The candidates are excluded when SGR J1935+2154 is eclipsed by Earth;
  2. 2.
    The candidates with the angle of the three most significant detectors relative to the source position (R.A. = 293.73°, decl. = 21.90°, Israel et al., 2016) should be less than $60^{\circ}$ (with at least two detectors);
  3. 3.
    The solar flares, the particle events measured by CPD observation, and the Terrestrial gamma-ray flash (TGF, Briggs et al., 2013) and Terrestrial electron beam (TEB, Xiong et al., 2012) featured by being harder ($\sim$ MeV) and shorter ($\sim$ 1 ms) than magnetar bursts are excluded;
  4. 4.
    Identified by other instruments (e.g. Fermi/GBM).
[…]
=-1cm
Table 1: Spectral templates used in the targeted search for X-ray bursts from SGR J1935+2154.
Spectrum Template alpha beta Epeak or kT (keV)
SoftBand -1.9 -3.7 70
CPL 0.25 28
OTTB 25
Blackbody 8
Powerlaw 2.4
Table 2: The proportion of different templates in X-ray bursts that GECAM-B/C observes from SGR J1935+2154.
Instrument Template Number Proportion
GECAM-B SoftBand 27 16.46%
CPL 39 23.78%
OTTB 57 34.75%
Blackbody 31 18.90%
Powerlaw 10 6.09%
GECAM-C SoftBand 7 7.21%
CPL 35 36.08%
OTTB 24 24.74%
Blackbody 19 19.58%
Powerlaw 12 12.37%
All bursts SoftBand 34 12.84%
CPL 74 17.15%
OTTB 81 23.04%
Blackbody 50 15.12%
Powerlaw 24 7.85%
2.2 Search Results
A total of 164 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a fast radio burst FRB 20221014 (Wang et al., 2022b). It is worth to note that about one third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM (blocked by the Earth or turned off during the SAA passage).
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to targeted search. The CPL, OTTB, and BB account for a proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai et al., 2021). As shown in Fig. 3, Using a Powerlaw template to search magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing much difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel (¿40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to found using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
[…]
Table 3: SGR J1935+2154 burst list observed by GECAM-B from January 2021 to December 2022.
Burst ID Trigger Time Template R.A. Decl. Err Sig.aaSignificance obtained by search pipeline. $T_{\mathrm{bb}}$bbBurst duration derived by a Bayesian block method. $T_{\mathrm{st}}$ccBurst start time relative to the trigger time. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBMddCalculated with the Fermi/GBM position history file (poshist) using the GBM data tool (Goldstein et al., 2022). GECAM-C
1 2021-01-02T03:15:43.330 OTTB 277.91 4.04 4.66 7.35 4.73 -2.58 Yes No
2 2021-01-06T05:11:16.390 BB 276.55 3.39 5.85 16.33 18.63 -8.80 Yes No
3 2021-01-24T00:00:59.050 CPL 290.82 4.82 15.64 11.82 13.39 -10.20 Yes No
4 2021-01-25T03:09:50.100 PL 290.44 3.23 11.80 7.30 130.96 -84.19 Yes No
5 2021-01-26T18:59:44.750 BB 297.66 25.20 8.51 8.35 38.18 -7.44 Yes No
6 2021-01-27T06:50:20.750 CPL 292.43 21.57 1.36 21.78 122.33 -36.92 No No
7 2021-01-29T17:51:00.800 PL 303.28 18.16 10.76 7.59 269.71 -170.50 Yes No
8 2021-01-30T08:39:53.840 PL 297.79 28.21 5.09 6.76 147.77 -28.85 Yes No
9 2021-01-30T08:48:23.580 OTTB 292.49 28.45 10.66 6.53 6.72 -3.53 Yes No
10 2021-01-30T10:35:35.120 BB 286.80 40.23 6.82 10.03 19.55 0.62 Yes No
11 2021-01-30T17:40:54.750 OTTB 292.67 22.92 1.35 7.27 121.16 -0.74 Yes No
12 2021-01-30T21:01:22.865 OTTB 286.55 23.59 19.83 6.37 72.04 -5.88 Yes No
13 2021-02-11T13:43:16.760 BB 275.98 4.12 10.73 7.64 28.50 -24.53 Yes No
14 2021-02-16T22:20:39.600 CPL 290.02 13.60 0.85 10.11 301.80 17.46 Yes No
15 2021-03-03T06:15:26.625 BB 277.83 3.80 3.76 7.06 26.09 -16.07 Yes No
16 2021-07-07T00:33:31.640 OTTB 292.75 23.87 1.68 13.88 124.74 10.48 Yes No
17 2021-07-08T00:18:18.560 SoftBand 278.35 23.77 24.26 6.90 302.37 -25.77 Yes No
18 2021-07-10T05:05:41.415 CPL 299.37 16.61 4.98 6.69 53.12 -5.03 No No
19 2021-07-12T04:32:39.600 OTTB 289.92 26.53 4.98 11.62 22.23 -32.91 No No
20 2021-07-12T22:12:58.100 BB 291.62 16.81 4.23 6.73 51.78 -68.85 No No
21 2021-08-24T07:53:15.060 OTTB 281.61 2.78 11.03 7.88 9.04 -7.03 Yes No
22 2021-09-09T20:58:35.600 BB 309.61 29.12 10.20 9.86 28.62 -39.05 No No
23 2021-09-09T21:07:12.150 CPL 295.39 23.87 3.57 7.32 55.13 10.39 No No
24 2021-09-10T01:04:33.500 OTTB 291.10 4.21 12.06 8.36 13.16 116.38 Yes No
25 2021-09-10T02:07:56.700 OTTB 288.57 18.71 9.07 6.66 71.03 -48.93 No No
26 2021-09-10T02:08:28.800 OTTB 291.44 27.38 4.75 6.64 145.70 -54.09 No No
27 2021-09-10T03:22:40.550 CPL 295.26 21.50 3.48 12.71 377.13 -19.89 No No
28 2021-09-10T03:24:47.150 OTTB 297.37 16.63 2.97 9.91 48.37 -9.69 No No
29 2021-09-10T03:42:45.750 CPL 280.88 19.63 10.79 9.63 52.04 -55.59 No No
30 2021-09-10T05:05:03.350 OTTB 302.51 16.83 5.93 10.22 302.41 -111.45 No No
31 2021-09-10T05:35:55.500 OTTB 302.09 24.12 5.78 10.21 21.73 -17.50 Yes No
32 2021-09-11T05:32:38.620 BB 292.32 26.42 2.39 7.70 221.50 -17.22 Yes No
33 2021-09-11T16:35:46.500 BB 296.83 9.68 10.67 6.80 42.40 -69.29 Yes No
34 2021-09-11T16:39:21.000 OTTB 299.09 7.43 7.63 7.82 53.30 -86.96 Yes No
35 2021-09-11T16:50:03.850 BB 293.68 27.82 3.20 9.78 39.63 -25.12 Yes No
36 2021-09-11T17:01:10.800 CPL 284.93 30.88 0.86 7.67 1349.53 -1050.55 Yes No
37 2021-09-11T17:04:29.800 BB 289.81 20.21 10.56 7.28 16.48 -49.93 Yes No
38 2021-09-11T17:10:48.750 OTTB 300.98 14.80 4.52 6.95 6.83 194.27 Yes No
39 2021-09-11T18:02:13.500 OTTB 309.94 5.57 10.59 7.66 38.60 -37.01 Yes No
40 2021-09-11T18:04:46.350 CPL 297.80 25.61 6.46 6.57 61.00 -28.29 Yes No
41 2021-09-11T18:54:36.050 BB 296.30 21.81 7.54 27.16 37.25 -20.13 Yes No
42 2021-09-11T19:43:28.000 OTTB 297.52 19.48 4.57 13.98 243.94 -83.05 Yes No
43 2021-09-11T19:46:50.050 OTTB 297.55 19.50 5.33 10.77 77.06 -84.11 Yes No
44 2021-09-11T20:13:40.550 BB 292.05 22.33 3.39 9.69 152.69 -63.17 Yes No
45 2021-09-11T20:22:59.050 CPL 292.04 22.33 5.15 7.13 180.72 -21.73 Yes No
46 2021-09-11T20:33:14.550 PL 281.74 36.29 6.24 6.39 87.57 -105.27 No No
47 2021-09-11T21:07:28.350 CPL 292.10 19.98 15.28 7.12 48.17 -183.59 Yes No
48 2021-09-11T22:51:41.600 CPL 297.54 19.47 9.35 17.29 116.25 -29.25 Yes No
49 2021-09-12T00:34:37.450 CPL 295.13 24.27 3.19 7.41 347.42 -41.07 Yes No
50 2021-09-12T00:45:49.400 BB 297.81 24.43 4.28 7.68 32.26 -14.45 Yes No
51 2021-09-12T05:14:07.950 CPL 298.65 17.10 5.87 13.87 110.85 -124.44 Yes No
52 2021-09-12T05:44:17.050 OTTB 306.96 11.58 15.77 7.23 80.14 -37.80 Yes No
53 2021-09-12T16:26:08.150 OTTB 308.99 17.73 7.83 8.80 66.08 -76.48 Yes No
54 2021-09-12T22:16:36.200 PL 287.91 29.81 10.18 10.08 8.48 -13.39 No No
55 2021-09-13T00:27:25.200 OTTB 297.19 16.71 9.10 8.78 252.68 -31.43 Yes No
56 2021-09-13T19:51:33.350 CPL 294.58 22.60 4.40 6.75 161.79 -183.78 Yes No
57 2021-09-14T11:10:36.250 CPL 294.53 22.56 3.91 19.75 94.94 -41.84 Yes No
58 2021-09-14T14:15:42.900 BB 293.45 13.94 6.45 9.10 13.05 -18.35 Yes No
59 2021-09-14T23:21:58.500 OTTB 293.08 24.79 15.42 6.48 131.71 -100.90 No No
60 2021-09-14T23:26:34.050 CPL 298.30 27.84 4.91 8.41 30.67 -25.93 No No
61 2021-09-15T02:35:47.350 CPL 299.49 2.24 12.11 7.90 46.36 -56.07 No No
62 2021-09-15T02:39:25.700 OTTB 302.11 21.02 6.28 9.06 66.16 -56.21 No No
63 2021-09-15T15:32:56.050 CPL 306.03 21.53 11.12 6.49 109.28 -97.54 Yes No
64 2021-09-17T12:52:37.800 CPL 295.59 19.10 11.61 8.54 32.38 -54.13 Yes No
65 2021-09-17T13:58:25.100 SoftBand 307.02 24.67 29.66 5.82 173.25 -191.07 Yes No
66 2021-09-18T22:58:52.150 OTTB 305.69 36.83 10.05 6.48 86.02 -89.76 Yes No
67 2021-09-22T02:39:10.200 CPL 295.06 22.95 10.30 8.96 114.19 -36.91 No No
68 2021-09-22T20:12:16.500 OTTB 286.20 25.10 0.87 8.78 170.69 -31.95 No No
69 2021-09-29T23:41:12.245 BB 292.03 24.98 11.59 15.86 18.85 -5.05 No No
70 2021-09-30T01:31:06.165 BB 292.04 22.52 7.96 10.14 50.53 -13.05 Yes No
71 2021-10-01T00:04:04.340 BB 300.29 25.34 10.11 15.07 29.50 -6.45 Yes No
72 2021-10-07T11:57:07.700 CPL 295.51 23.25 4.32 7.93 310.13 -129.76 No No
73 2021-10-11T07:41:20.085 BB 281.55 3.28 3.11 11.69 11.68 -6.89 Yes No
74 2021-11-01T23:13:41.950 OTTB 291.66 23.42 9.52 6.74 26.91 -41.10 No No
75 2022-01-04T04:32:11.200 BB 291.23 16.98 8.33 10.50 35.71 -55.83 Yes No
76 2022-01-05T06:01:31.450 BB 297.80 20.99 4.91 6.56 75.39 -96.74 Yes No
77 2022-01-05T07:06:40.800 CPL 294.95 22.51 4.25 17.59 126.75 -52.60 Yes No
78 2022-01-06T02:36:14.100 OTTB 298.33 22.47 1.66 28.11 51.67 -54.52 Yes No
79 2022-01-08T14:41:46.900 SoftBand 296.90 18.66 7.73 5.10 76.18 -31.01 Yes No
80 2022-01-09T07:39:10.700 CPL 298.56 24.55 5.29 7.16 35.87 -18.96 Yes No
81 2022-01-10T02:57:16.825 BB 297.95 20.19 3.40 27.78 84.87 -5.71 Yes No
82 2022-01-10T06:52:40.500 CPL 289.20 19.47 4.23 6.46 63.07 -34.89 No No
83 2022-01-11T08:58:35.450 BB 296.97 20.80 3.07 10.91 171.34 -110.70 Yes No
84 2022-01-12T01:03:46.900 OTTB 288.32 19.51 4.73 7.18 116.26 -111.12 Yes No
85 2022-01-12T02:19:22.200 CPL 291.72 21.95 4.59 7.05 392.65 -142.87 Yes No
86 2022-01-12T05:42:51.650 CPL 295.89 23.01 5.18 8.65 198.73 -173.05 Yes No
87 2022-01-12T08:39:25.450 OTTB 291.94 23.59 0.44 7.86 979.51 -183.09 Yes No
88 2022-01-12T17:57:08.500 OTTB 290.11 27.80 4.14 7.14 147.71 -143.36 Yes No
89 2022-01-13T19:36:08.600 CPL 291.78 22.50 6.73 7.08 48.18 -86.71 Yes No
90 2022-01-13T20:06:58.760 OTTB 292.59 30.85 10.13 8.45 37.51 -11.38 Yes No
91 2022-01-13T20:14:58.600 OTTB 300.47 19.00 3.47 9.81 459.77 -453.04 Yes No
92 2022-01-13T21:41:17.900 CPL 298.29 25.00 5.74 9.69 46.64 -52.24 Yes No
93 2022-01-14T19:42:08.050 CPL 294.18 20.61 2.09 21.27 234.33 805.01 Yes No
94 2022-01-14T19:45:08.100 CPL 291.46 18.19 5.96 6.71 27.85 -37.33 Yes No
95 2022-01-14T19:56:52.700 OTTB 296.11 22.79 0.69 57.43 382.90 -265.83 Yes No
96 2022-01-14T20:06:07.400 OTTB 291.24 27.15 6.01 6.34 36.48 -62.88 No No
97 2022-01-14T20:07:03.050 OTTB 291.78 22.66 1.10 8.92 315.07 -6.13 No No
98 2022-01-14T20:12:45.300 BB 292.07 15.64 7.45 7.38 511.42 -525.26 No No
99 2022-01-14T20:15:54.400 OTTB 290.77 22.18 0.96 7.23 279.46 -131.00 No No
100 2022-01-14T20:21:05.150 OTTB 293.51 23.17 0.15 8.20 1552.57 -95.36 No No
101 2022-01-14T20:22:53.955 BB 297.11 21.96 5.04 13.85 360.17 -3.72 No No
102 2022-01-14T20:23:35.400 BB 297.49 23.27 1.41 7.26 197.65 -153.85 No No
103 2022-01-14T20:26:50.300 BB 289.22 19.27 6.06 7.29 17.96 -29.98 No No
104 2022-01-14T20:29:07.250 OTTB 296.99 20.38 0.86 23.31 629.51 -127.86 No No
105 2022-01-14T20:31:49.900 OTTB 292.19 15.06 5.05 8.57 197.92 -148.02 No No
106 2022-01-15T09:26:39.900 SoftBand 294.31 19.69 4.34 19.15 38.68 -36.41 Yes No
107 2022-01-15T13:52:26.050 CPL 294.53 19.68 5.77 12.82 72.31 -42.00 No No
108 2022-01-15T16:31:14.900 CPL 299.83 24.81 6.13 7.14 270.39 -263.97 Yes No
109 2022-01-15T17:21:59.300 OTTB 295.64 20.99 1.28 14.43 246.44 -28.19 Yes No
110 2022-01-16T10:48:37.650 CPL 296.68 22.09 3.38 21.66 281.15 -51.90 Yes No
111 2022-01-17T01:27:12.720 OTTB 297.06 16.94 8.47 8.03 88.85 -35.25 Yes No
112 2022-01-17T01:39:37.300 OTTB 303.85 24.20 11.21 7.34 136.90 -92.61 Yes No
113 2022-01-20T18:52:48.950 OTTB 291.96 26.72 6.02 6.75 31.61 -41.89 No No
114 2022-01-23T20:06:38.750 CPL 297.12 20.14 2.10 16.17 357.03 -8.50 No No
115 2022-01-24T02:10:55.050 CPL 293.10 23.40 2.88 8.79 251.47 -47.28 No No
116 2022-02-02T14:52:28.170 BB 296.53 19.32 9.85 7.04 30.11 -8.93 No No
117 2022-02-06T05:15:45.635 CPL 286.83 23.23 4.08 10.00 43.00 -7.87 Yes No
118 2022-05-21T23:32:58.800 BB 290.97 25.93 5.26 8.76 39.62 -28.50 Yes No
119 2022-05-23T23:29:22.850 OTTB 285.67 22.26 10.49 7.19 90.88 -22.11 No No
120 2022-05-24T17:10:18.550 CPL 297.09 21.45 4.20 14.34 30.38 -36.51 No No
121 2022-10-11T14:11:36.850 SoftBand 308.83 22.03 3.41 18.79 35.22 -49.06 Yes Yes
122 2022-10-12T12:47:04.400 SoftBand 290.23 21.22 2.88 7.55 113.52 -15.72 Yes Yes
123 2022-10-12T14:22:47.000 SoftBand 302.37 14.62 9.13 15.70 22.80 -24.19 Yes Yes
124 2022-10-12T14:40:50.450 SoftBand 299.91 24.47 8.50 16.04 41.38 -22.76 No No
125 2022-10-12T14:43:19.950 SoftBand 295.94 27.86 2.59 44.72 184.74 -60.42 No No
126 2022-10-12T14:44:45.600 SoftBand 294.85 22.46 0.48 16.26 848.21 -53.68 No No
127 2022-10-12T14:54:37.950 SoftBand 301.62 18.24 14.63 7.86 9.94 -51.20 No Yes
128 2022-10-12T15:42:31.250 SoftBand 278.03 34.16 1.39 7.60 395.20 -20.92 Yes Yes
129 2022-10-12T15:45:10.150 SoftBand 281.00 26.32 0.83 6.88 393.21 -196.87 Yes Yes
130 2022-10-12T22:34:56.900 SoftBand 299.32 5.40 9.37 7.50 29.84 -43.19 No No
131 2022-10-13T00:00:53.450 SoftBand 291.38 20.32 1.10 8.41 348.81 -92.94 Yes No
132 2022-10-14T06:47:24.100 SoftBand 299.64 22.19 5.12 13.70 242.99 -150.79 No Yes
133 2022-10-14T07:12:28.800 SoftBand 300.00 19.45 3.44 89.16 214.46 -74.52 Yes Yes
134 2022-10-14T07:24:42.150 SoftBand 298.97 25.74 10.83 12.28 76.77 -24.94 Yes Yes
135 2022-10-14T11:27:32.750 SoftBand 298.24 22.50 3.87 31.96 25.27 -28.72 No Yes
136 2022-10-14T13:21:36.400 SoftBand 291.14 24.79 4.09 28.45 164.56 -62.66 Yes Yes
137 2022-10-14T17:04:16.650 CPL 278.57 2.86 5.35 7.92 33.32 -36.02 Yes No
138 2022-10-14T19:21:39.100 SoftBand 296.04 21.09 4.38 6.88 185.32 -56.10 No Yes
139 2022-10-16T21:14:58.450 SoftBand 305.40 21.07 10.12 14.14 27.16 -55.42 Yes No
140 2022-10-16T21:20:52.250 SoftBand 297.92 20.53 1.70 29.83 103.84 -17.82 Yes No
141 2022-10-17T15:40:10.400 BB 277.95 25.72 4.88 22.52 146.21 -21.22 No Yes
142 2022-10-17T16:24:53.900 BB 293.60 23.57 10.62 15.46 49.42 -36.06 Yes Yes
143 2022-10-22T01:41:18.050 SoftBand 298.31 24.43 7.37 17.47 19.45 -44.49 No No
144 2022-10-22T04:44:52.700 SoftBand 293.12 22.71 4.41 26.81 38.10 -24.84 No Yes
145 2022-11-08T19:48:49.900 PL 282.55 27.31 16.97 8.63 23.47 -38.93 Yes Yes
146 2022-11-09T14:33:10.200 OTTB 291.93 23.64 3.99 10.70 376.85 -97.67 Yes Yes
147 2022-11-09T16:06:08.620 OTTB 296.91 23.11 1.48 7.24 2505.18 -44.20 Yes Yes
148 2022-11-09T17:46:18.375 PL 290.42 24.43 7.66 7.80 145.80 -25.70 Yes Yes
149 2022-11-09T17:47:59.220 OTTB 294.18 22.41 4.03 7.18 320.31 -23.75 Yes Yes
150 2022-11-09T17:56:52.715 OTTB 288.28 16.94 7.89 9.88 36.62 -22.52 Yes No
151 2022-11-09T18:25:56.380 OTTB 281.78 26.43 4.74 9.08 93.27 -12.64 No Yes
152 2022-11-10T06:39:03.075 SoftBand 312.13 22.87 10.13 7.87 129.89 -78.53 Yes No
153 2022-11-10T09:54:23.100 OTTB 290.04 18.06 4.66 19.78 92.53 -20.96 No No
154 2022-11-10T15:15:13.355 OTTB 286.61 20.67 10.87 9.75 23.17 -9.98 Yes Yes
155 2022-11-10T18:10:56.200 PL 303.35 25.14 4.04 9.54 25.16 -23.64 No Yes
156 2022-11-11T13:11:49.125 OTTB 292.34 29.21 3.54 8.27 10.56 -2.46 No Yes
157 2022-11-12T00:21:33.900 OTTB 292.05 23.64 3.53 8.31 146.83 -60.74 No Yes
158 2022-11-12T03:32:17.705 OTTB 292.32 19.92 5.00 11.18 11.13 -5.18 No Yes
159 2022-11-15T17:14:34.860 OTTB 290.93 23.59 5.55 14.86 164.34 -6.85 No Yes
160 2022-11-19T21:53:33.050 PL 289.43 32.60 14.68 7.10 35.58 -48.63 Yes Yes
161 2022-11-20T11:25:04.850 SoftBand 304.95 22.04 4.77 6.86 283.82 -85.17 Yes No
162 2022-12-13T06:57:10.650 OTTB 300.02 20.96 2.51 12.81 133.42 -68.79 Yes No
163 2022-12-16T20:20:17.855 PL 281.43 2.58 1.95 9.30 79.04 -5.84 Yes Yes
164 2022-12-25T00:21:47.700 OTTB 278.16 2.57 2.08 7.13 135.67 -2.18 No Yes
[…]
Table 4: SGR J1935+2154 burst list observed by GECAM-C from October to December 2022.
Burst ID Trigger Time Template R.A. Decl. Err Sig.aaSignificance obtained by search pipeline. $T_{\mathrm{bb}}$bbBurst duration derived by a Bayesian block method. $T_{\mathrm{st}}$ccBurst start time relative to the trigger time. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBMddCalculated with the Fermi/GBM position history file (poshist) using the GBM data tool (Goldstein et al., 2022). GECAM-B
1 2022-10-09T10:32:49.680 OTTB 311.94 10.14 6.50 6.99 6.73 -2.26 Yes Yes
2 2022-10-11T10:41:16.655 CPL 276.75 5.08 3.76 10.39 51.70 -6.90 Yes No
3 2022-10-12T12:47:04.425 OTTB 296.47 30.54 4.89 16.11 496.77 -299.79 Yes Yes
4 2022-10-12T15:06:51.200 SoftBand 309.59 38.81 8.85 7.86 39.92 -82.27 No Yes
5 2022-10-12T15:07:36.710 BB 295.79 23.31 5.47 6.55 651.24 -8.37 No Yes
6 2022-10-12T15:08:48.575 OTTB 299.27 19.45 4.48 8.51 779.10 -8.11 No Yes
7 2022-10-12T15:11:19.095 BB 297.75 20.67 3.99 10.13 297.69 -163.07 No Yes
8 2022-10-12T15:14:04.100 BB 294.14 22.16 2.87 12.30 724.92 -10.46 Yes No
9 2022-10-12T15:14:34.525 CPL 294.18 20.63 3.49 6.59 320.39 -64.49 Yes No
10 2022-10-12T15:14:51.325 BB 297.43 26.44 9.57 9.96 44.96 -41.58 Yes No
11 2022-10-12T15:15:58.465 PL 289.17 23.23 9.50 9.73 37.52 -6.34 Yes No
12 2022-10-12T15:16:22.975 CPL 289.97 18.96 5.85 8.17 1219.09 -36.12 Yes No
13 2022-10-12T15:16:45.925 OTTB 299.67 16.11 9.47 7.03 7.25 -7.07 Yes No
14 2022-10-12T15:18:51.805 CPL 286.25 17.67 5.14 17.48 98.53 -21.56 Yes No
15 2022-10-12T15:20:30.930 BB 293.34 22.84 1.98 10.83 2661.75 -25.13 Yes No
16 2022-10-12T15:21:39.825 CPL 299.55 14.37 7.04 10.68 11.79 -6.85 Yes No
17 2022-10-12T15:22:25.140 CPL 290.27 22.46 4.15 11.02 1117.16 -8.40 Yes No
18 2022-10-12T15:49:13.545 PL 304.99 42.33 5.47 8.69 439.68 -208.56 Yes Yes
19 2022-10-12T18:33:43.480 CPL 303.91 23.43 0.84 11.33 92.65 -9.66 Yes No
20 2022-10-13T02:13:21.465 OTTB 280.81 10.28 3.26 17.04 148.57 -37.56 No Yes
21 2022-10-13T13:59:10.900 BB 290.36 19.93 11.36 11.40 6.45 -9.10 Yes No
22 2022-10-13T15:30:31.525 CPL 293.56 23.82 4.23 8.28 830.16 -38.84 Yes No
23 2022-10-13T22:41:29.435 CPL 299.20 30.35 0.32 22.25 501.80 -82.95 No Yes
24 2022-10-13T22:55:29.320 BB 293.50 12.22 9.99 9.89 27.75 -11.30 Yes Yes
25 2022-10-14T02:27:47.140 BB 293.15 22.85 4.02 7.97 54.90 -2.92 Yes Yes
26 2022-10-14T05:42:15.425 OTTB 293.30 34.32 13.08 8.35 111.34 -26.66 Yes Yes
27 2022-10-14T05:47:52.600 OTTB 282.28 24.40 6.14 7.10 317.33 -65.42 Yes Yes
28 2022-10-14T06:47:24.045 OTTB 287.80 16.05 8.99 6.58 265.49 -102.01 No Yes
29 2022-10-14T07:12:28.740 CPL 291.66 18.70 2.81 7.06 215.53 -13.67 Yes Yes
30 2022-10-14T07:17:36.155 OTTB 289.57 16.41 5.80 7.06 306.75 -18.63 Yes Yes
31 2022-10-14T09:56:04.850 CPL 308.30 5.13 5.09 7.84 149.19 -96.41 No Yes
32 2022-10-14T10:28:39.825 OTTB 290.95 26.03 6.72 7.36 139.10 -33.50 Yes Yes
33 2022-10-14T11:21:05.325 CPL 284.30 24.38 8.98 10.75 136.81 -44.76 No Yes
34 2022-10-14T11:27:32.730 CPL 293.39 19.48 3.52 30.99 44.95 -13.14 No Yes
35 2022-10-14T11:58:40.230 OTTB 284.20 15.58 5.05 9.75 85.89 -10.25 Yes Yes
36 2022-10-14T12:47:44.045 BB 301.50 27.49 3.25 29.16 52.81 -4.77 No Yes
37 2022-10-14T13:20:31.720 PL 291.35 24.31 7.06 8.90 244.57 -44.63 Yes Yes
38 2022-10-14T13:21:36.375 CPL 288.70 23.05 2.93 9.55 691.62 -16.21 Yes Yes
39 2022-10-14T13:38:25.400 OTTB 311.32 12.79 10.29 8.01 55.88 -67.71 Yes Yes
40 2022-10-14T17:35:01.675 SoftBand 278.66 31.91 9.25 7.67 47.79 -38.67 No Yes
41 2022-10-14T17:47:59.175 SoftBand 280.58 31.65 1.46 7.20 486.30 -130.42 No Yes
42 2022-10-14T19:21:39.020 OTTB 302.72 18.95 5.94 8.03 321.19 -10.00 No Yes
43 2022-10-14T20:48:47.425 SoftBand 286.06 37.31 11.61 6.97 254.97 -102.61 No Yes
44 2022-10-15T02:13:54.545 SoftBand 288.25 4.63 1.26 11.28 128.58 -17.94 Yes Yes
45 2022-10-16T10:43:17.080 BB 282.62 19.36 8.50 8.81 22.04 -3.94 No Yes
46 2022-10-17T04:43:27.720 CPL 276.95 9.82 7.64 7.45 16.37 -10.85 Yes No
47 2022-10-17T14:17:40.650 OTTB 291.35 21.34 10.91 10.52 47.56 -15.97 Yes No
48 2022-10-17T15:14:07.600 PL 280.53 12.30 9.83 6.61 268.73 -221.31 No Yes
49 2022-10-17T15:53:30.340 BB 279.93 16.88 14.97 7.50 13.36 -12.99 Yes No
50 2022-10-17T16:35:38.085 CPL 281.66 12.46 8.73 9.82 48.81 -26.24 Yes Yes
51 2022-10-18T13:57:25.965 SoftBand 286.87 2.99 10.27 7.23 4.82 -10.47 Yes Yes
52 2022-10-18T16:22:09.650 CPL 300.25 25.07 8.56 11.35 295.70 -5.49 Yes No
53 2022-10-22T04:44:52.700 CPL 300.97 29.09 3.36 21.45 86.99 -4.82 No Yes
54 2022-10-22T16:30:16.000 PL 301.02 40.42 4.85 9.18 109.45 -134.88 Yes No
55 2022-10-25T18:42:31.610 BB 277.10 4.59 12.13 7.22 14.00 -6.36 No No
56 2022-10-31T23:08:59.180 CPL 277.14 15.63 2.91 10.49 14.27 -4.90 Yes Yes
57 2022-11-09T09:05:47.515 CPL 291.45 14.72 3.44 8.74 298.00 -10.95 No No
58 2022-11-09T14:33:10.135 CPL 288.10 19.74 3.98 26.27 432.66 -7.65 Yes Yes
59 2022-11-09T15:29:02.435 CPL 286.99 18.24 3.06 8.52 502.87 -1.47 Yes No
60 2022-11-09T15:58:26.975 OTTB 288.62 26.59 15.04 7.32 71.15 -69.01 Yes Yes
61 2022-11-09T16:06:08.595 CPL 290.43 21.32 2.45 6.07 2537.76 -3.71 Yes Yes
62 2022-11-09T16:49:33.240 CPL 294.07 32.15 10.32 14.52 643.86 -14.48 No Yes
63 2022-11-09T16:50:28.385 OTTB 292.75 26.98 18.72 7.35 39.40 -27.64 No Yes
64 2022-11-09T16:51:37.330 PL 280.08 26.93 8.11 13.33 31.15 -5.87 No No
65 2022-11-09T16:52:01.750 OTTB 295.82 27.61 5.45 6.92 33.43 223.88 No No
66 2022-11-09T16:52:26.585 OTTB 278.83 23.98 9.29 6.68 111.64 -29.51 No No
67 2022-11-09T16:54:17.565 CPL 295.70 23.93 3.80 15.72 359.39 -13.91 No No
68 2022-11-09T16:55:20.665 BB 287.46 29.18 13.97 7.39 30.20 -10.96 No No
69 2022-11-09T16:57:30.950 BB 278.25 15.69 17.94 7.45 69.34 -13.18 No No
70 2022-11-09T16:59:56.580 CPL 290.30 20.83 3.57 19.85 100.90 -3.42 No No
71 2022-11-09T17:05:04.670 CPL 286.18 22.81 4.97 12.97 480.98 -1.80 Yes No
72 2022-11-09T17:21:41.635 OTTB 293.34 21.22 2.71 11.74 169.43 -8.33 Yes No
73 2022-11-09T17:35:49.100 OTTB 303.74 24.36 4.51 10.84 331.01 -29.99 Yes Yes
74 2022-11-09T17:46:18.390 BB 299.15 21.10 3.62 7.27 144.09 0.58 Yes Yes
75 2022-11-09T17:47:59.245 CPL 290.50 21.31 2.98 13.92 307.82 -7.27 Yes Yes
76 2022-11-09T18:25:56.375 CPL 289.59 26.79 2.63 17.78 96.94 -10.60 No Yes
77 2022-11-09T19:07:34.000 CPL 293.20 13.03 16.49 8.16 22.47 -9.50 Yes Yes
78 2022-11-09T19:08:04.250 SoftBand 276.47 9.37 2.08 8.84 97.96 -51.83 Yes Yes
79 2022-11-09T20:09:32.965 BB 289.05 27.24 6.08 10.98 69.20 -15.85 No No
80 2022-11-09T20:45:50.955 OTTB 284.97 21.56 4.86 8.45 223.92 -12.21 Yes Yes
81 2022-11-10T09:03:33.235 CPL 302.00 22.68 6.94 10.81 99.04 -11.88 Yes No
82 2022-11-10T14:01:58.135 OTTB 284.30 13.73 7.35 6.67 25.53 -5.10 Yes No
83 2022-11-10T18:10:56.205 OTTB 289.45 23.58 5.15 19.23 40.50 -5.87 No Yes
84 2022-11-10T23:43:45.995 CPL 285.79 13.09 4.22 12.58 43.11 -6.08 Yes No
85 2022-11-13T02:56:52.190 CPL 279.47 13.13 4.74 13.81 41.79 -7.96 No No
86 2022-11-14T11:09:19.010 BB 285.86 14.33 5.62 13.49 53.19 -4.31 No No
87 2022-11-15T17:14:34.865 CPL 294.84 21.22 2.91 10.28 181.28 -10.70 No Yes
88 2022-11-19T18:42:35.280 OTTB 284.95 12.68 1.96 15.21 6.65 -4.78 Yes Yes
89 2022-11-23T09:35:02.435 PL 294.73 7.75 14.95 7.37 4.97 -5.21 Yes Yes
90 2022-11-23T19:17:25.100 PL 311.40 24.56 8.94 6.53 2.82 -17.85 Yes Yes
91 2022-11-27T00:38:15.935 PL 310.60 35.64 10.30 7.98 8.41 -8.30 Yes Yes
92 2022-11-30T01:05:01.125 PL 299.69 17.49 4.36 7.01 8.99 -13.51 No Yes
93 2022-12-01T17:31:06.735 BB 276.56 38.99 4.63 7.40 9.78 -9.50 Yes Yes
94 2022-12-05T16:51:36.925 PL 306.92 13.65 7.23 6.94 1.94 -10.05 Yes No
95 2022-12-19T08:49:13.535 BB 284.86 20.70 17.26 7.52 6.97 -7.84 Yes No
96 2022-12-20T13:09:17.025 PL 310.28 18.36 8.90 7.38 6.70 -9.98 Yes Yes
97 2022-12-21T03:26:06.710 CPL 289.30 12.99 4.84 7.96 8.65 -8.95 Yes Yes
Refer to captionRefer to caption
Figure 1: The targeted search location (TL) of bursts. The red star marks the accurate position of SGR J1935+2154 (R.A. = 293.73°, decl. = 21.90°).
Refer to captionRefer to caption
Figure 2: The angle separation between the targeted search location (TL) of burst and true position of SGR J1935+2154.
Refer to captionRefer to caption
Figure 3: The significance of burst v.s angle separation between the targeted search location (TL) and true position of SGR J1935+2154.
Figure 4: Representative examples of GECAM-detected bursts of SGR J1935+2154. For each burst, light curves are shown for high gain (HG, 6–300 keV). The background of each burst is shown by the red dotted line. The green vertical dotted line and blue dotted line are assessed by the Bayesian block (BB) algorithm.
3 Catalog Analysis
As shown above, we show light curves of 9 bursts in Fig. 4 to display the variety of these bursts. The trigger time of each burst is listed in Tables 3 and 4. SGR J1935+2154 bursts consist of a single pulse (such as ID: 122, 138, and 161 of GECAM-B; ID: 3 and 42 of GECAM-C), multiple pulses (ID: 27 of GECAM-B; ID: 59 of GECAM-C), and intermediate flares (ID: 126 of GECAM-B; ID: 23 of GECAM-C). The X-ray burst associated with the fast radio burst FRB 20221014 is ID: 138 of GECAM-B, and ID: 42 of GECAM-C.
3.1 Burst Activity
The daily burst rate is defined as $N/P$, where $N$ is the observed burst number per day and $P$ represents the percentage of the effective observation time in a full day, excluding the time intervals when the Earth blocked SGR J1935+2154, or when GECAM passes through the South Atlantic Anomaly (SAA) area, during the instrument’s startup period.
We calculate the effective observation time  [^3] GECAM-B orbits at low latitudes ($<30^{\circ}N/S$), while GECAM-C could cover nearly $90^{\circ}N/S$. We only calculate the observation time of GECAM-C at low latitudes in this paper since the background undergoes too violent changes to detect magnetar bursts when GECAM-C is across high latitudes. of GECAM-B/C individually and estimate the daily burst rate of GECAM-B/C, respectively (see Fig. 5). Because the effective detectable energy range of GECAM-B changed in the HG channel (¿40 keV after October 2022), weaker bursts may not be detected compared to GECAM-C observations. Hence, given that the burst activity of GECAM-B is lower than that of GECAM-C (¿15 keV), the daily burst rate estimated using GECAM-C data may be more reliable for assessing burst activity compared to using GECAM-B data after October 2022. A detailed study on the evolution of the effective detectable energy range change of GECAM-B will be reported in Paper II.
Figure 5: The burst history of SGR J1935+2154 observed by GECAM-B/C.
3.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data from GECAM-B (before October 2022) and GECAM-C (after October 2022) with the Lomb-Scargle Periodogram for searching periodicity (Lomb, 1976; Scargle, 1982; VanderPlas, 2018). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (False Alarm Probability: 0.001). The peak around the 50 day are likely caused by observation windows or bursts gap (Xie et al., 2022).
To uncover the active period of SGR J1935+2154, the start time of all bursts are folded into different phases using period folding method (CHIME/FRB Collaboration et al., 2020b; Zhang et al., 2021; Xie et al., 2022),
$\phi=\frac{T-T_{0}}{P}-\mathrm{floor}\biggl{(}\frac{T-T_{0}}{P}\biggr{)},$ (6)
where $\phi$ is the folded phase, $T$ is the burst start time, $T_{0}$ is the phase start point (MJD: 59215; UTC: 2021-01-01), $P$ is a given period, and the floor is a function that returns the floor of the input number. As shown in Fig 7, the duty circle of SGR J1935+2154 bursts is 80%. The error could be estimated to be 20 days with
$\sigma=\frac{PW_{\mathrm{active}}}{T_{\mathrm{span}}},$ (7)
where $P$ is the period and $W_{\mathrm{active}}$ is the number of active days. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is $134.63\pm 20$ days, consistent with previous work (Xie et al., 2022, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
Figure 6: Lomb-Scargle periodogram of X-ray bursts from SGR J1935+2154 observed by GECAM. The vertical red dotted line indicate the peak of periodogram which is 134.63 day.
Figure 7: Phase-folded burst rate according to the 134.63 day period with MJD 59215 (2021-01-01) referenced as phase 0. The active window (i.e. from $\phi=0.7$ to $\phi=1.5$) is denoted with red dotted lines.
3.3 Burst Duration
As previous studies (Lin et al., 2013), the magnetar burst duration $T_{\mathrm{bb}}$ is calculated with the Bayesian blocks method (Scargle et al., 2013). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s are treated as background since the spin period of SGR J1935+2154 is just approximately 3.24 s. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance ($S/\sqrt{B}$, where $S$ is the net counts and $B$ is the background counts) larger than $5\sigma$ is considered as part of the burst region. As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple pulses (e.g. burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the entire burst duration is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
The duration ($T_{\mathrm{bb}}$) of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We conduct the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin et al., 2020b, a; Cai et al., 2022a; Rehan & Ibrahim, 2023, 2024).
[…]
=-1cm
Table 5: The mean of burst duration distribution in each episode obtained by fitting to a lognormal function.
Instrument Episode Mean Statisticaaone-sample Kolmogorov-Smirnov test statistic p-value
GECAM-B 1st episode $45.50_{-32.56}^{+114.46}$ 0.18 0.64
2nd episode $65.77_{-43.12}^{+125.22}$ 0.06 0.98
3rd episode $120.78_{-79.38}^{+231.59}$ 0.11 0.57
4th episode $86.30_{-60.65}^{+204.10}$ 0.11 0.58
GECAM-C 4th episode $83.37_{-66.35}^{+324.95}$ 0.07 0.69
Figure 8: The burst duration of 14 X-ray bursts observed by both GECAM-B and GECAM-C.
Figure 9: The left panel is the burst duration of each burst observed by GECAM-B/C. The right panel is the burst duration distributions of GECAM-B/C. The horizontal dotted lines show the best-fit lognormal function with a mean value listed in Table 5.
3.4 Burst Waiting Time
The waiting time between the successive bursts that fall within an uninterrupted observation time interval is defined as
$\Delta{t}=t_{\mathrm{i}}-t_{\mathrm{i-1}}$ (8)
where $t_{i}$ is the burst start time of the $i-$th burst. As shown in Fig. 10, there are 45 waiting times of GECAM-B and 21 waiting times of GECAM-C during continuous observation time intervals.
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and a mean value of $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C. When comparing to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution mean value of GECAM-C is smaller than that of GECAM-B.
Figure 10: The left panel is the waiting time (time interval between two successive bursts observed by GECAM-B/C) of the X-ray bursts v.s their occurring times, where the time of FRB 20221014 is represented by the green dashed line. The right panel is the distribution of waiting time. The horizontal dotted lines represent the best-fit lognormal function with a mean of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C, respectively.
3.5 Burst Hardness Ratio
The net counts of the source is estimated as
$N_{\mathrm{i}}=C_{\mathrm{i}}-B_{\mathrm{i}},$ (9)
where $C_{\mathrm{i}}$ and $B_{\mathrm{i}}$ represent the total counts and background counts in the burst duration, respectively. The hardness ratio is the ratio of the net counts in different energy ranges.
Since changes in the effective detectable energy range of GECAM-B over the two years, we measure the hardness ratio of all bursts within the same effective range of 50-200/40-50 keV for GECAM-B. The hardness ratio of GECAM-C is calculated in the range of 30-200/15-30 keV. The hardness ratio of the four active episodes for GECAM-B data yields mean values of 2.70, 2.30, 1.87, and 1.48, respectively. It presents that the evolution of the burst hardness tends to be softer across the observation time, as shown in Fig. 11. The burst hardness is more concentrated after October 2022 when FRB 20221014 happened. More detailed spectral properties and evolution will be reported in Paper II.
Figure 11: The evolution of hardness ratio of each burst in 50-200/40-50 keV for GECAM-B and 30-200/15-30 keV for GECAM-C.
4 Summary and Discussions
In this paper, we perform a targeted sub-threshold search GECAM daily EVT data for SGR J1935+2154 bursts during the year of 2021 and 2022. We design 5 spectral templates to target search and find 164 X-ray bursts from GECAM-B and 97 bursts from GECAM-C (refer to Tables 3 and 4). Among the templates, the CPL, OTTB, and BB cover a higher proportion of them compared to others (refer to Table 2), providing more accuracy in location (see Fig. 2). Therefore, a simple power-law function and exponential cutoff function with softer parameter settings might be preferred for searching magnetar X-ray bursts.
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, resulting in the most likely period of $134.63\pm 20$ days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into 4 active episodes (see Fig. 5). The period of $134.63\pm 20$ days with 80% duty circle is well consistent with previous study (Xie et al., 2022, evaluated from July 2014 to January 2022). Unlike other study periods ($\sim 231$ days evaluated from 2014 to 2020 or $\sim 238$ days evaluated from July 2014 to October 2021 in Grossan, 2021; Zou et al., 2021, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy ranges when both satellites detect the same burst. The waiting time also conforms to a lognormal function with mean values of $338.84_{-212.95}^{+573.16}$ seconds and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-B/C, respectively (refer to Fig. 10). Because the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). The hardness ratio of X-ray bursts tends to become softer and more concentrated over the course of the observation time spanning these two years (see Fig. 11).
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity and exhibit a softer spectrum over the two years. Similar to FRB 200428, which occurred during the most active period (April-May 2020) preceding 2020, the FRB 20221014 are detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular, magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017). And the amounts of energy released by the bursts and outbursts of Galactic magnetars, from SGRs or AXPs, indicate that X-ray bursts probably correspond to sudden and local (not global) catastrophic events. In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang & Zhang, 2021; Xie et al., 2024).
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550300, XDA30050000), the National Key R&D Program of China (2021YFA0718500), the National SKA program of China (2020SKA0120300), the National Natural Science Foundation of China (Grant No. 12393811, 12173038, 12303045, 12273042), the Strategic Priority Research Program on Space Science of the Chinese Academy of Sciences (Grant No. XDA15360102, XDA15360300, XDA15052700), the Natural Science Foundation of Hebei Province (No. A2023205020), and the Science Research Program of Dezhou University (2024xjrc142).