### v5

# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

Abstract
Magnetar is a neutron star with an ultrahigh magnetic field ($\sim 10^{14}-10^{15}$ G). The magnetar SGR J1935+2154 is not only one of the most active magnetars detected so far, but also the unique confirmed source of fast radio bursts (FRBs). Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM) is dedicated to monitor gamma-ray transients all over the sky, including magnetar short bursts. Here we report the GECAM observations of the burst activity of SGR J1935+2154 from January 2021 to December 2022, which results in a unique and valuable data set for this important magnetar. With a targeted search of GECAM data, 159 bursts from SGR J1935+2154 are detected by GECAM-B while 97 bursts by GECAM-C, including the X-ray burst associated with a bright radio burst. We find that both the burst duration and the waiting time between two successive bursts follow lognormal distributions. The period of burst activity is $134\pm 20$ days, thus the burst activity could be generally divided into four active episodes over these two years. Interestingly, the hardness ratio of X-ray bursts tends to be softer during these two years, especially during the active episode with radio bursts detected.
magnetars - soft gamma-ray repeaters: general - methods: data analysis - techniques
1 Introduction
Magnetars are highly magnetized isolated neutron stars (Duncan & Thompson, 1992; van Kerkwijk et al., 1995; Kouveliotou et al., 1998; Banas et al., 1997; Kaspi & Beloborodov, 2017), and are characterized by their slow rotation period (P212similar-toabsent212$\sim$ 2-12∼ 2 - 12 s), rapidly spinning down ($\dot{P}\sim 10^{-13}-10^{-11}$ $\mathrm{s\cdot s^{-1}}$), and relatively young age (typically about 1000 yr). We currently know about 30 magnetars (Olausen & Kaspi, 2014) [^1] http://www.physics.mcgill.ca/~pulsar/magnetar/main.html. Magnetars are manifested themselves by persistent emission and emit bursts/flares simultaneously in the X-/gamma-ray band (Esposito et al., 2021).
Based on their luminosity and duration, magnetar bursts can be divided into three classes (Woods & Thompson, 2006): Short-duration burst, which consists of single or multiple peaks, is the most typical magnetar burst with a burst duration $\sim$0.01 $\mathrm{s}$ and the luminosity around $10^{38}\sim 10^{40}$ $\mathrm{erg\cdot s^{-1}}$; Intermediate burst is a brighter magnetar burst with a duration longer than a short-duration burst ($>1$ $\mathrm{s}$) and peak luminosity around $10^{41}\sim 10^{43}$ $\mathrm{erg\cdot s^{-1}}$; Giant flare, the rarest and the most powerful energetic burst is characterized by a significantly higher luminosity than a typical magnetar burst and a light curve with a hard initial spike and rapidly decaying tail modulated at the spin period of the magnetar.
Magnetar SGR J1935+2154 was first discovered and located in the Milky Way Galaxy by the Swift Burst Alert Telescope (BAT) in 2014 July (Stamatikos et al., 2014), while it is worth noting that Xie et al. (2022) found several burst candidates in the gamma-ray burst monitor of Fermi gamma-ray Space Telescope (Fermi/GBM, Meegan et al., 2009) before this time. Follow-up observations carried out between 2014 July and 2015 March with Chandra and XMM-Newton allowed the measurement of its spin period and spin-down rate, $P\sim 3.24\ \mathrm{s}$ and $\dot{P}\sim 1.43\times 10^{-11}\ \mathrm{s\cdot s^{-1}}$, respectively. This indicates the estimated magnetic field at the pole is $B\sim 2.2\times 10^{14}\ \mathrm{G}$ (Israel et al., 2016). It has experienced multiple outbursts from 2014 to 2022 (Younes et al., 2017; Lin et al., 2020b, a; Cai et al., 2022a, b; Xie et al., 2022; Rehan & Ibrahim, 2023, 2024). April 2020 was recognized as a month of intense bursting activity for SGR J1935+2154, during which burst forests were observed (Kaneko et al., 2021; Younes et al., 2020). Moreover, the source entered a new outburst (Borghese et al., 2022). Among these bursts there is an X-ray burst (Li et al., 2021; Mereghetti et al., 2020; Ridnaia et al., 2021; Tavani et al., 2020) associated with a fast radio burst, FRB 200428 (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020a). Additional coincident radio and X-ray bursts were observed during the active episode of 2022 (Li et al., 2021; Xiao et al., 2022a; Zhang et al., 2023; Wang et al., 2024, 2022a; Li et al., 2022).
In this work, we carry out a comprehensive targeted search on X-ray bursts of SGR J1935+2154 using GECAM observations data from January 2021 to December 2022. As in Paper I of this series work for GECAM observations of SGR J1935+2154, we focus on the burst search, identification, and temporal analyses. In Paper II (in prep.), we will give a detailed time-integrated spectral analysis of these bursts. The current paper is organized as follows: In Section 2, we present the observations, the search process, and report the burst catalog. In Section 3, we present the catalog analysis method and the results. Finally, a summary and discussions are given in Section 4.
Please note that the error reported in this work is for 1 sigma, if not stated otherwise.
2 Observation and Burst Search
As of the writing of this paper, GECAM constellation is composed of four instruments, GECAM-A/B (were launched in December 2020, Xiao et al., 2022b), GECAM-C (i.e. SATech-01/HEBS, launched in July 2022, Zhang et al., 2023), and GECAM-D (i.e. DRO/GTM, launched in March 2024, Wang et al., 2024). These four satellites are all-sky monitors with a large field of view that could monitor various gamma-ray transients except for the area blocked by Earth. Both GECAM-A and GECAM-B feature a dome-shaped array of 25 gamma-ray detectors (GRDs) and 8 charged particle detectors (CPDs), while GECAM-C has 12 GRDs and 2 CPDs. All GECAM-B’s GRDs and most GECAM-C’s GRDs operate in two readout channels: high gain (HG, 6–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The event-by-event (EVT) data from GECAM’s GRD detectors is used for further analysis. The GECAM achieves a high time resolution of 0.1 $\mathrm{\mu s}$, with a dead time of 4 $\mathrm{\mu s}$ for normal events and exceeding 69 $\mathrm{\mu s}$ for overflow events (Liu et al., 2021). We note that in this work only GECAM-B/C data are used because of power supply issues that GECAM-A has not been able to observe the source yet (Li et al., 2022).
2.1 Search and Identification
Magnetar SGR J1935+2154 experienced multiple active X-ray burst episodes after the launch of GECAM-A/B, and an active X-ray burst episode after the launch of GECAM-C (Rehan & Ibrahim, 2023, 2024; Wood & Fermi-GBM Team, 2022). A search pipeline developed by Cai et al. (2025) for gamma-ray bursts (GRBs) is used to do a blind search on GECAM daily observation data. In addition to other GRBs, 82 X-ray bursts from SGR J1935+2154 were also found by GECAM-B (2021-01 to 2022-01, Xie et al., 2022) [^2] The 94 bursts shown in Xie et al. (2022) are the preliminary blind searched results that are reported in the Gamma-ray Coordinates Network (GCN) by the burst advocate (BA). We identify 82 bursts from SGR J1935+2154 using the targeted search in this paper..
The search pipeline is implemented using a ground sensitive coherent search method that features a log-likelihood ratio (hereafter LR) comparing expected model counts to observed counts in each channel of each detector (Blackburn et al., 2015; Cai et al., 2025),
$\mathcal{L}=\sum_{i=1}^{j}[\ln\frac{\sigma_{n_{\mathrm{i}}}}{\sigma_{d_{\mathrm{i}}}}+\frac{\widetilde{d_{\mathrm{i}}^{2}}}{2\sigma_{n_{\mathrm{i}}}^{2}}-\frac{(\widetilde{d_{\mathrm{i}}}-r_{\mathrm{i}}s)^{2}}{2\sigma_{d_{\mathrm{i}}}^{2}}],$ (1)
where $i$ is the number of data sets in each channel and detector, $j$ is the total number of detectors and channels, $\widetilde{d_{\mathrm{i}}^{2}}$ is the background-subtracted counts and $\sigma_{d_{\mathrm{i}}}$ is the error of the expected counts (background+source), $n_{\mathrm{i}}$ is the estimated background and $\sigma_{n_{\mathrm{i}}}$ is the error of the background counts, $r_{\mathrm{i}}$ and $s$ represent the instrument response and the expected source amplitude (could estimated by a spectral model template), respectively. Therefore, the LR could be calculated with the response factors over the all-sky for a blind search or a certain area for a targeted search. Then, the continuous time-bins with the higher LR (higher than the LR threshold) could be considered as one signal (trigger/burst). The maximum LR gives the most possible spatial location of the source signal. The pipeline will mark the first time-bin, of which LR is higher than the LR threshold, as a trigger time. In this paper, the LR threshold is set to 20 and the search timescale (time-bin) is set to 10 ms.
To unveil potential bursts that are relatively weak and unable to trigger the instrument with a normal threshold (typically $5\sigma$ as implemented in flight software), we conduct a targeted sub-threshold search for the area around the position of SGR J1935+2154 (R.A. and Decl. ±$\pm 20^{\circ}$) on GECAM daily EVT data [^3] https://gecamweb.ihep.ac.cn/dailydatadownload.jhtml from 2021-01-01 to 2022-12-31 using the pipeline developed by Cai et al. (2025). Since X-ray bursts from magnetars tend to be softer than general GRBs, it is essential to employ appropriate spectral model templates instead of default Band templates for the search. Based on previous studies on magnetar burst spectral properties (Lin et al., 2020b, a; Rehan & Ibrahim, 2023, 2024), we utilize the following spectral model templates: the soft Band function (default softer template for GRB searching), CPL (a power law with high-energy exponential cutoff), OTTB (optically-thin thermal bremsstrahlung), Blackbody, and Powerlaw. The parameter setting is listed in Table 1 which is similar to Table 1 in Xie et al. (2022). The details of each template are presented below.
The Band function:
$A(E)\sim\begin{cases}E^{\alpha}\exp\left[-\frac{(2+\alpha)E}{E_{\mathrm{p}}}\right],&\text{if }E<\frac{(\alpha-\beta)E_{\mathrm{p}}}{2+\alpha},\\ \\ E^{\beta}\exp(\beta-\alpha)\left[\frac{(\alpha-\beta)E_{\mathrm{p}}}{(2+\alpha)}\right]^{\alpha-\beta},&\text{otherwise,}\end{cases}$ (2)
where $E_{\mathrm{p}}$, in units of keV, is the peak energy of the $\nu F_{\mathrm{\nu}}$ spectrum, $\alpha$/$\beta$ are the low-/high-energy photon indices.
The CPL function:
$A(E)\sim E^{-\alpha}\exp\left[-\frac{(2-\alpha)E}{E_{\mathrm{p}}}\right],$ (3)
where $E_{\mathrm{p}}$, in units of keV, is the peak energy of the $\nu F_{\mathrm{\nu}}$ spectrum, $\alpha$ are the photon index.
The OTTB function:
$A(E)\sim E^{-1}\exp\left(-\frac{E}{kT}\right),$ (4)
where $kT$ is the electron energy, in units of keV.
The Blackbody function:
$A(E)\sim\frac{E^{2}}{\exp(E/kT)-1},$ (5)
where $kT$ is the temperature, in units of keV.
The Powerlaw function:
$A(E)\sim E^{-\alpha},$ (6)
where $\alpha$ are the photon index.
Thousands of triggers (candidate bursts) have been found through the targeted search process. We perform a refined analysis for each candidate burst by carefully removing various false triggers or other burst events using the GRB data analysis tool, GECAMTools [^4] https://github.com/zhangpeng-sci/GECAMTools-Public. SGR J1935+2154 burst identified from the candidates should adhere to the following criteria (same as Xie et al., 2022):
  1. 1.
    Using the GECAM position history file (posatt) to retrieve the position of satellite and the pointing of detector. The candidates are excluded when SGR J1935+2154 is occulted by Earth;
  2. 2.
    The burst location is within the error circle (at 3 $\sigma$) of the source position (SGR J1935+2154, R.A. = 293.73°, Decl. = 21.90°, Israel et al., 2016);
  3. 3.
    The solar flares, the particle events measured by CPD observations, and the Terrestrial gamma-ray flash (TGF, Briggs et al., 2013) and Terrestrial electron beam (TEB, Xiong et al., 2012) featured by being harder ($\sim$ MeV) and shorter ($\sim$ 1 ms) than magnetar bursts are excluded;
  4. 4.
    The pipeline will calculate the significance [^5] In this paper, the significance is defined as $S/\sqrt{B}$, where $S$ is the source (background-subtracted) counts and $B$ is the estimated background counts. of each detector. At least two most significant detectors should have an angle relative to the source position of less than $60^{\circ}$ or maybe a false trigger (e.g., particle events);
  5. 5.
    Identified by other instruments (e.g., Fermi/GBM).
[…]
=-1cm
Table 1: Spectral templates used in the targeted search for X-ray bursts from SGR J1935+2154.
Spectrum Template alpha beta Epeak or kT (keV)
SoftBand -1.9 -3.7 70
CPL 0.25 28
OTTB 25
Blackbody 8
Powerlaw 2.4
Table 2: The proportion of different templates in X-ray bursts that GECAM-B/C observes from SGR J1935+2154.
Instrument Template Number Proportion
GECAM-B SoftBand 27 16.98%
CPL 39 24.52%
OTTB 55 34.59%
Blackbody 29 18.23%
Powerlaw 9 5.66%
GECAM-C SoftBand 7 7.21%
CPL 35 36.08%
OTTB 24 24.74%
Blackbody 19 19.58%
Powerlaw 12 12.37%
All bursts SoftBand 34 24.19%
CPL 74 60.60%
OTTB 79 59.33%
Blackbody 48 37.81%
Powerlaw 21 18.03%
2.2 Search Results
A total of 159 bursts for GECAM-B and 97 bursts for GECAM-C are identified and listed in Tables 3 and 4. Among them, there are 14 bursts that both GECAM-B and GECAM-C could observe, and a short X-ray burst (2022-10-14T19:21:39.130 UTC) observed by GECAM-B&C is associated with a bright radio burst FRB 20221014 (Dong & Chime/Frb Collaboration, 2022; Wang et al., 2022b; Frederiks et al., 2022; Maan et al., 2022; Giri et al., 2023) [^6] We marker this radio burst as FRB 20221014 in this paper since the luminosity is several orders higher than the pulsed emission from SGR J1935+2154 (Zhu et al., 2023) although still a few orders lower than that of general FRBs.. It is worth noting that about one-third of GECAM bursts (i.e. 60 bursts of GECAM-B and 36 bursts of GECAM-C) are invisible to Fermi/GBM [^7] Using the Fermi/GBM position history file (poshist) and the GBM data tool (Goldstein et al., 2022) to determine if SGR J1935+2154 is visible to Fermi/GBM or occulted by the Earth at these burst times..
The targeted search location (TL) of each burst is plotted in Fig 1. We summarize the proportion of different templates in Table 2 and Fig. 2 to assess each template’s merit for magnetar bursts to the targeted search. The CPL, OTTB, and BB account for a proportion of them than other templates. Note that even a burst could be found by every template, the pipeline only outputs the best template (e.g., the one with the maximum log-likelihood ratio, as referenced in Cai et al., 2021). As shown in Fig. 3, using a Powerlaw template to search for a magnetar burst prefers a weak burst with low significance. Most of the SoftBand templates are given by GECAM-B data after October 2022, showing a much difference compared to those from GECAM-C. This disparity may be attributed to the effective detectable energy range of GECAM-B changed in the HG channel ($>$40 keV after October 2022), potentially leading to bursts with more high energy photons being detected without softer photons, making them easier to found using the SoftBand compared to other templates. Therefore, in general, it is recommended to use a template with a simple power-law function and exponential cutoff function, such as CPL, OTTB, or Blackbody, to coherently search for and locate magnetar bursts in hard X-ray observations.
[…]
Table 3: SGR J1935+2154 burst list observed by GECAM-B from January 2021 to December 2022.
Burst ID Trigger Time Template R.A. Decl. Err Sig. $T_{\mathrm{bb}}$aaBurst duration derived by the Bayesian blocks algorithm. $T_{\mathrm{st}}$bbBurst start time relative to the trigger time using the Bayesian blocks algorithm. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBM GECAM-C
1 2021-01-02T03:15:43.330 OTTB 277.91 4.04 4.66 7.35 4.73 -2.58 Yes No
2 2021-01-06T05:11:16.390 BB 276.55 3.39 5.85 16.33 18.63 -8.80 Yes No
3 2021-01-24T00:00:59.050 CPL 290.82 4.82 15.64 11.82 13.39 -10.20 Yes No
4 2021-01-25T03:09:50.100 PL 290.44 3.23 11.80 7.30 130.96 -84.19 Yes No
5 2021-01-26T18:59:44.750 BB 297.66 25.20 8.51 8.35 38.18 -7.44 Yes No
6 2021-01-27T06:50:20.750 CPL 292.43 21.57 1.36 21.78 122.33 -36.92 No No
7 2021-01-29T17:51:00.800 PL 303.28 18.16 10.76 7.59 269.71 -170.50 Yes No
8 2021-01-30T08:39:53.840 PL 297.79 28.21 5.09 6.76 147.77 -28.85 Yes No
9 2021-01-30T08:48:23.580 OTTB 292.49 28.45 10.66 6.53 6.72 -3.53 Yes No
10 2021-01-30T10:35:35.120 BB 286.80 40.23 6.82 10.03 19.55 0.62 Yes No
11 2021-01-30T17:40:54.750 OTTB 292.67 22.92 1.35 7.27 121.16 -0.74 Yes No
12 2021-01-30T21:01:22.865 OTTB 286.55 23.59 19.83 6.37 72.04 -5.88 Yes No
13 2021-02-11T13:43:16.760 BB 275.98 4.12 10.73 7.64 28.50 -24.53 Yes No
14 2021-02-16T22:20:39.600 CPL 290.02 13.60 0.85 10.11 301.80 17.46 Yes No
15 2021-07-07T00:33:31.640 OTTB 292.75 23.87 1.68 13.88 124.74 10.48 Yes No
16 2021-07-08T00:18:18.560 SoftBand 278.35 23.77 24.26 6.90 302.37 -25.77 Yes No
17 2021-07-10T05:05:41.415 CPL 299.37 16.61 4.98 6.69 53.12 -5.03 No No
18 2021-07-12T04:32:39.600 OTTB 289.92 26.53 4.98 11.62 22.23 -32.91 No No
19 2021-07-12T22:12:58.100 BB 291.62 16.81 4.23 6.73 51.78 -68.85 No No
20 2021-09-09T20:58:35.600 BB 309.61 29.12 10.20 9.86 28.62 -39.05 No No
21 2021-09-09T21:07:12.150 CPL 295.39 23.87 3.57 7.32 55.13 10.39 No No
22 2021-09-10T01:04:33.500 OTTB 291.10 4.21 12.06 8.36 13.16 116.38 Yes No
23 2021-09-10T02:07:56.700 OTTB 288.57 18.71 9.07 6.66 71.03 -48.93 No No
24 2021-09-10T02:08:28.800 OTTB 291.44 27.38 4.75 6.64 145.70 -54.09 No No
25 2021-09-10T03:22:40.550 CPL 295.26 21.50 3.48 12.71 377.13 -19.89 No No
26 2021-09-10T03:24:47.150 OTTB 297.37 16.63 2.97 9.91 48.37 -9.69 No No
27 2021-09-10T03:42:45.750 CPL 280.88 19.63 10.79 9.63 52.04 -55.59 No No
28 2021-09-10T05:05:03.350 OTTB 302.51 16.83 5.93 10.22 302.41 -111.45 No No
29 2021-09-10T05:35:55.500 OTTB 302.09 24.12 5.78 10.21 21.73 -17.50 Yes No
30 2021-09-11T05:32:38.620 BB 292.32 26.42 2.39 7.70 221.50 -17.22 Yes No
31 2021-09-11T16:35:46.500 BB 296.83 9.68 10.67 6.80 42.40 -69.29 Yes No
32 2021-09-11T16:39:21.000 OTTB 299.09 7.43 7.63 7.82 53.30 -86.96 Yes No
33 2021-09-11T16:50:03.850 BB 293.68 27.82 3.20 9.78 39.63 -25.12 Yes No
34 2021-09-11T17:01:10.800 CPL 284.93 30.88 0.86 7.67 1349.53 -1050.55 Yes No
35 2021-09-11T17:04:29.800 BB 289.81 20.21 10.56 7.28 16.48 -49.93 Yes No
36 2021-09-11T17:10:48.750 OTTB 300.98 14.80 4.52 6.95 6.83 194.27 Yes No
37 2021-09-11T18:02:13.500 OTTB 309.94 5.57 10.59 7.66 38.60 -37.01 Yes No
38 2021-09-11T18:04:46.350 CPL 297.80 25.61 6.46 6.57 61.00 -28.29 Yes No
39 2021-09-11T18:54:36.050 BB 296.30 21.81 7.54 27.16 37.25 -20.13 Yes No
40 2021-09-11T19:43:28.000 OTTB 297.52 19.48 4.57 13.98 243.94 -83.05 Yes No
41 2021-09-11T19:46:50.050 OTTB 297.55 19.50 5.33 10.77 77.06 -84.11 Yes No
42 2021-09-11T20:13:40.550 BB 292.05 22.33 3.39 9.69 152.69 -63.17 Yes No
43 2021-09-11T20:22:59.050 CPL 292.04 22.33 5.15 7.13 180.72 -21.73 Yes No
44 2021-09-11T20:33:14.550 PL 281.74 36.29 6.24 6.39 87.57 -105.27 No No
45 2021-09-11T21:07:28.350 CPL 292.10 19.98 15.28 7.12 48.17 -183.59 Yes No
46 2021-09-11T22:51:41.600 CPL 297.54 19.47 9.35 17.29 116.25 -29.25 Yes No
47 2021-09-12T00:34:37.450 CPL 295.13 24.27 3.19 7.41 347.42 -41.07 Yes No
48 2021-09-12T00:45:49.400 BB 297.81 24.43 4.28 7.68 32.26 -14.45 Yes No
49 2021-09-12T05:14:07.950 CPL 298.65 17.10 5.87 13.87 110.85 -124.44 Yes No
50 2021-09-12T05:44:17.050 OTTB 306.96 11.58 15.77 7.23 80.14 -37.80 Yes No
51 2021-09-12T16:26:08.150 OTTB 308.99 17.73 7.83 8.80 66.08 -76.48 Yes No
52 2021-09-12T22:16:36.200 PL 287.91 29.81 10.18 10.08 8.48 -13.39 No No
53 2021-09-13T00:27:25.200 OTTB 297.19 16.71 9.10 8.78 252.68 -31.43 Yes No
54 2021-09-13T19:51:33.350 CPL 294.58 22.60 4.40 6.75 161.79 -183.78 Yes No
55 2021-09-14T11:10:36.250 CPL 294.53 22.56 3.91 19.75 94.94 -41.84 Yes No
56 2021-09-14T14:15:42.900 BB 293.45 13.94 6.45 9.10 13.05 -18.35 Yes No
57 2021-09-14T23:21:58.500 OTTB 293.08 24.79 15.42 6.48 131.71 -100.90 No No
58 2021-09-14T23:26:34.050 CPL 298.30 27.84 4.91 8.41 30.67 -25.93 No No
59 2021-09-15T02:35:47.350 CPL 299.49 2.24 12.11 7.90 46.36 -56.07 No No
60 2021-09-15T02:39:25.700 OTTB 302.11 21.02 6.28 9.06 66.16 -56.21 No No
61 2021-09-15T15:32:56.050 CPL 306.03 21.53 11.12 6.49 109.28 -97.54 Yes No
62 2021-09-17T12:52:37.800 CPL 295.59 19.10 11.61 8.54 32.38 -54.13 Yes No
63 2021-09-17T13:58:25.100 SoftBand 307.02 24.67 29.66 5.82 173.25 -191.07 Yes No
64 2021-09-18T22:58:52.150 OTTB 305.69 36.83 10.05 6.48 86.02 -89.76 Yes No
65 2021-09-22T02:39:10.200 CPL 295.06 22.95 10.30 8.96 114.19 -36.91 No No
66 2021-09-22T20:12:16.500 OTTB 286.20 25.10 0.87 8.78 170.69 -31.95 No No
67 2021-09-29T23:41:12.245 BB 292.03 24.98 11.59 15.86 18.85 -5.05 No No
68 2021-09-30T01:31:06.165 BB 292.04 22.52 7.96 10.14 50.53 -13.05 Yes No
69 2021-10-01T00:04:04.340 BB 300.29 25.34 10.11 15.07 29.50 -6.45 Yes No
70 2021-10-07T11:57:07.700 CPL 295.51 23.25 4.32 7.93 310.13 -129.76 No No
71 2021-11-01T23:13:41.950 OTTB 291.66 23.42 9.52 6.74 26.91 -41.10 No No
72 2022-01-04T04:32:11.200 BB 291.23 16.98 8.33 10.50 35.71 -55.83 Yes No
73 2022-01-05T06:01:31.450 BB 297.80 20.99 4.91 6.56 75.39 -96.74 Yes No
74 2022-01-05T07:06:40.800 CPL 294.95 22.51 4.25 17.59 126.75 -52.60 Yes No
75 2022-01-06T02:36:14.100 OTTB 298.33 22.47 1.66 28.11 51.67 -54.52 Yes No
76 2022-01-08T14:41:46.900 SoftBand 296.90 18.66 7.73 5.10 76.18 -31.01 Yes No
77 2022-01-09T07:39:10.700 CPL 298.56 24.55 5.29 7.16 35.87 -18.96 Yes No
78 2022-01-10T02:57:16.825 BB 297.95 20.19 3.40 27.78 84.87 -5.71 Yes No
79 2022-01-10T06:52:40.500 CPL 289.20 19.47 4.23 6.46 63.07 -34.89 No No
80 2022-01-11T08:58:35.450 BB 296.97 20.80 3.07 10.91 171.34 -110.70 Yes No
81 2022-01-12T01:03:46.900 OTTB 288.32 19.51 4.73 7.18 116.26 -111.12 Yes No
82 2022-01-12T02:19:22.200 CPL 291.72 21.95 4.59 7.05 392.65 -142.87 Yes No
83 2022-01-12T05:42:51.650 CPL 295.89 23.01 5.18 8.65 198.73 -173.05 Yes No
84 2022-01-12T08:39:25.450 OTTB 291.94 23.59 0.44 7.86 979.51 -183.09 Yes No
85 2022-01-12T17:57:08.500 OTTB 290.11 27.80 4.14 7.14 147.71 -143.36 Yes No
86 2022-01-13T19:36:08.600 CPL 291.78 22.50 6.73 7.08 48.18 -86.71 Yes No
87 2022-01-13T20:06:58.760 OTTB 292.59 30.85 10.13 8.45 37.51 -11.38 Yes No
88 2022-01-13T20:14:58.600 OTTB 300.47 19.00 3.47 9.81 459.77 -453.04 Yes No
89 2022-01-13T21:41:17.900 CPL 298.29 25.00 5.74 9.69 46.64 -52.24 Yes No
90 2022-01-14T19:42:08.050 CPL 294.18 20.61 2.09 21.27 234.33 805.01 Yes No
91 2022-01-14T19:45:08.100 CPL 291.46 18.19 5.96 6.71 27.85 -37.33 Yes No
92 2022-01-14T19:56:52.700 OTTB 296.11 22.79 0.69 57.43 382.90 -265.83 Yes No
93 2022-01-14T20:06:07.400 OTTB 291.24 27.15 6.01 6.34 36.48 -62.88 No No
94 2022-01-14T20:07:03.050 OTTB 291.78 22.66 1.10 8.92 315.07 -6.13 No No
95 2022-01-14T20:12:45.300 BB 292.07 15.64 7.45 7.38 511.42 -525.26 No No
96 2022-01-14T20:15:54.400 OTTB 290.77 22.18 0.96 7.23 279.46 -131.00 No No
97 2022-01-14T20:21:05.150 OTTB 293.51 23.17 0.15 8.20 1552.57 -95.36 No No
98 2022-01-14T20:22:53.955 BB 297.11 21.96 5.04 13.85 360.17 -3.72 No No
99 2022-01-14T20:23:35.400 BB 297.49 23.27 1.41 7.26 197.65 -153.85 No No
100 2022-01-14T20:26:50.300 BB 289.22 19.27 6.06 7.29 17.96 -29.98 No No
101 2022-01-14T20:29:07.250 OTTB 296.99 20.38 0.86 23.31 629.51 -127.86 No No
102 2022-01-14T20:31:49.900 OTTB 292.19 15.06 5.05 8.57 197.92 -148.02 No No
103 2022-01-15T09:26:39.900 SoftBand 294.31 19.69 4.34 19.15 38.68 -36.41 Yes No
104 2022-01-15T13:52:26.050 CPL 294.53 19.68 5.77 12.82 72.31 -42.00 No No
105 2022-01-15T16:31:14.900 CPL 299.83 24.81 6.13 7.14 270.39 -263.97 Yes No
106 2022-01-15T17:21:59.300 OTTB 295.64 20.99 1.28 14.43 246.44 -28.19 Yes No
107 2022-01-16T10:48:37.650 CPL 296.68 22.09 3.38 21.66 281.15 -51.90 Yes No
108 2022-01-17T01:27:12.720 OTTB 297.06 16.94 8.47 8.03 88.85 -35.25 Yes No
109 2022-01-17T01:39:37.300 OTTB 303.85 24.20 11.21 7.34 136.90 -92.61 Yes No
110 2022-01-20T18:52:48.950 OTTB 291.96 26.72 6.02 6.75 31.61 -41.89 No No
111 2022-01-23T20:06:38.750 CPL 297.12 20.14 2.10 16.17 357.03 -8.50 No No
112 2022-01-24T02:10:55.050 CPL 293.10 23.40 2.88 8.79 251.47 -47.28 No No
113 2022-02-02T14:52:28.170 BB 296.53 19.32 9.85 7.04 30.11 -8.93 No No
114 2022-02-06T05:15:45.635 CPL 286.83 23.23 4.08 10.00 43.00 -7.87 Yes No
115 2022-05-21T23:32:58.800 BB 290.97 25.93 5.26 8.76 39.62 -28.50 Yes No
116 2022-05-23T23:29:22.850 OTTB 285.67 22.26 10.49 7.19 90.88 -22.11 No No
117 2022-05-24T17:10:18.550 CPL 297.09 21.45 4.20 14.34 30.38 -36.51 No No
118 2022-10-11T14:11:36.850 SoftBand 308.83 22.03 3.41 18.79 35.22 -49.06 Yes No
119 2022-10-12T12:47:04.400 SoftBand 290.23 21.22 2.88 7.55 113.52 -15.72 Yes No
120 2022-10-12T14:22:47.000 SoftBand 302.37 14.62 9.13 15.70 22.80 -24.19 Yes Yes
121 2022-10-12T14:40:50.450 SoftBand 299.91 24.47 8.50 16.04 41.38 -22.76 No Yes
122 2022-10-12T14:43:19.950 SoftBand 295.94 27.86 2.59 44.72 184.74 -60.42 No Yes
123 2022-10-12T14:44:45.600 SoftBand 294.85 22.46 0.48 16.26 848.21 -53.68 No No
124 2022-10-12T14:54:37.950 SoftBand 301.62 18.24 14.63 7.86 9.94 -51.20 No No
125 2022-10-12T15:42:31.250 SoftBand 278.03 34.16 1.39 7.60 395.20 -20.92 Yes No
126 2022-10-12T15:45:10.150 SoftBand 281.00 26.32 0.83 6.88 393.21 -196.87 Yes Yes
127 2022-10-12T22:34:56.900 SoftBand 299.32 5.40 9.37 7.50 29.84 -43.19 No Yes
128 2022-10-13T00:00:53.450 SoftBand 291.38 20.32 1.10 8.41 348.81 -92.94 Yes Yes
129 2022-10-14T06:47:24.100 SoftBand 299.64 22.19 5.12 13.70 242.99 -150.79 No No
130 2022-10-14T07:12:28.800 SoftBand 300.00 19.45 3.44 89.16 214.46 -74.52 Yes No
131 2022-10-14T07:24:42.150 SoftBand 298.97 25.74 10.83 12.28 76.77 -24.94 Yes Yes
132 2022-10-14T11:27:32.750 SoftBand 298.24 22.50 3.87 31.96 25.27 -28.72 No Yes
133 2022-10-14T13:21:36.400 SoftBand 291.14 24.79 4.09 28.45 164.56 -62.66 Yes Yes
134 2022-10-14T17:04:16.650 CPL 278.57 2.86 5.35 7.92 33.32 -36.02 Yes Yes
135 2022-10-14T19:21:39.100 SoftBand 296.04 21.09 4.38 6.88 185.32 -56.10 No Yes
136 2022-10-16T21:14:58.450 SoftBand 305.40 21.07 10.12 14.14 27.16 -55.42 Yes No
137 2022-10-16T21:20:52.250 SoftBand 297.92 20.53 1.70 29.83 103.84 -17.82 Yes Yes
138 2022-10-17T15:40:10.400 BB 277.95 25.72 4.88 22.52 146.21 -21.22 No No
139 2022-10-17T16:24:53.900 BB 293.60 23.57 10.62 15.46 49.42 -36.06 Yes No
140 2022-10-22T01:41:18.050 SoftBand 298.31 24.43 7.37 17.47 19.45 -44.49 No Yes
141 2022-10-22T04:44:52.700 SoftBand 293.12 22.71 4.41 26.81 38.10 -24.84 No Yes
142 2022-11-08T19:48:49.900 PL 282.55 27.31 16.97 8.63 23.47 -38.93 Yes No
143 2022-11-09T14:33:10.200 OTTB 291.93 23.64 3.99 10.70 376.85 -97.67 Yes Yes
144 2022-11-09T16:06:08.620 OTTB 296.91 23.11 1.48 7.24 2505.18 -44.20 Yes Yes
145 2022-11-09T17:46:18.375 PL 290.42 24.43 7.66 7.80 145.80 -25.70 Yes Yes
146 2022-11-09T17:47:59.220 OTTB 294.18 22.41 4.03 7.18 320.31 -23.75 Yes Yes
147 2022-11-09T17:56:52.715 OTTB 288.28 16.94 7.89 9.88 36.62 -22.52 Yes Yes
148 2022-11-09T18:25:56.380 OTTB 281.78 26.43 4.74 9.08 93.27 -12.64 No Yes
149 2022-11-10T06:39:03.075 SoftBand 312.13 22.87 10.13 7.87 129.89 -78.53 Yes No
150 2022-11-10T09:54:23.100 OTTB 290.04 18.06 4.66 19.78 92.53 -20.96 No Yes
151 2022-11-10T15:15:13.355 OTTB 286.61 20.67 10.87 9.75 23.17 -9.98 Yes No
152 2022-11-10T18:10:56.200 PL 303.35 25.14 4.04 9.54 25.16 -23.64 No No
153 2022-11-11T13:11:49.125 OTTB 292.34 29.21 3.54 8.27 10.56 -2.46 No Yes
154 2022-11-12T00:21:33.900 OTTB 292.05 23.64 3.53 8.31 146.83 -60.74 No Yes
155 2022-11-12T03:32:17.705 OTTB 292.32 19.92 5.00 11.18 11.13 -5.18 No Yes
156 2022-11-15T17:14:34.860 OTTB 290.93 23.59 5.55 14.86 164.34 -6.85 No Yes
157 2022-11-19T21:53:33.050 PL 289.43 32.60 14.68 7.10 35.58 -48.63 Yes Yes
158 2022-11-20T11:25:04.850 SoftBand 304.95 22.04 4.77 6.86 283.82 -85.17 Yes Yes
159 2022-12-13T06:57:10.650 OTTB 300.02 20.96 2.51 12.81 133.42 -68.79 Yes Yes
[…]
Table 4: SGR J1935+2154 burst list observed by GECAM-C from October to December 2022.
Burst ID Trigger Time Template R.A. Decl. Err Sig. $T_{\mathrm{bb}}$aaBurst duration derived by the Bayesian blocks algorithm. $T_{\mathrm{st}}$bbBurst start time relative to the trigger time using the Bayesian blocks algorithm. Visible to Visible to
(UTC) (deg) (deg) (deg) ($\sigma$) (ms) (ms) GBM GECAM-B
1 2022-10-09T10:32:49.680 OTTB 311.94 10.14 6.50 6.99 6.73 -2.26 Yes Yes
2 2022-10-11T10:41:16.655 CPL 276.75 5.08 3.76 10.39 51.70 -6.90 Yes No
3 2022-10-12T12:47:04.425 OTTB 296.47 30.54 4.89 16.11 496.77 -299.79 Yes Yes
4 2022-10-12T15:06:51.200 SoftBand 309.59 38.81 8.85 7.86 39.92 -82.27 No Yes
5 2022-10-12T15:07:36.710 BB 295.79 23.31 5.47 6.55 651.24 -8.37 No Yes
6 2022-10-12T15:08:48.575 OTTB 299.27 19.45 4.48 8.51 779.10 -8.11 No Yes
7 2022-10-12T15:11:19.095 BB 297.75 20.67 3.99 10.13 297.69 -163.07 No Yes
8 2022-10-12T15:14:04.100 BB 294.14 22.16 2.87 12.30 724.92 -10.46 Yes No
9 2022-10-12T15:14:34.525 CPL 294.18 20.63 3.49 6.59 320.39 -64.49 Yes No
10 2022-10-12T15:14:51.325 BB 297.43 26.44 9.57 9.96 44.96 -41.58 Yes No
11 2022-10-12T15:15:58.465 PL 289.17 23.23 9.50 9.73 37.52 -6.34 Yes No
12 2022-10-12T15:16:22.975 CPL 289.97 18.96 5.85 8.17 1219.09 -36.12 Yes No
13 2022-10-12T15:16:45.925 OTTB 299.67 16.11 9.47 7.03 7.25 -7.07 Yes No
14 2022-10-12T15:18:51.805 CPL 286.25 17.67 5.14 17.48 98.53 -21.56 Yes No
15 2022-10-12T15:20:30.930 BB 293.34 22.84 1.98 10.83 2661.75 -25.13 Yes No
16 2022-10-12T15:21:39.825 CPL 299.55 14.37 7.04 10.68 11.79 -6.85 Yes No
17 2022-10-12T15:22:25.140 CPL 290.27 22.46 4.15 11.02 1117.16 -8.40 Yes No
18 2022-10-12T15:49:13.545 PL 304.99 42.33 5.47 8.69 439.68 -208.56 Yes Yes
19 2022-10-12T18:33:43.480 CPL 303.91 23.43 0.84 11.33 92.65 -9.66 Yes No
20 2022-10-13T02:13:21.465 OTTB 280.81 10.28 3.26 17.04 148.57 -37.56 No Yes
21 2022-10-13T13:59:10.900 BB 290.36 19.93 11.36 11.40 6.45 -9.10 Yes No
22 2022-10-13T15:30:31.525 CPL 293.56 23.82 4.23 8.28 830.16 -38.84 Yes No
23 2022-10-13T22:41:29.435 CPL 299.20 30.35 0.32 22.25 501.80 -82.95 No Yes
24 2022-10-13T22:55:29.320 BB 293.50 12.22 9.99 9.89 27.75 -11.30 Yes Yes
25 2022-10-14T02:27:47.140 BB 293.15 22.85 4.02 7.97 54.90 -2.92 Yes Yes
26 2022-10-14T05:42:15.425 OTTB 293.30 34.32 13.08 8.35 111.34 -26.66 Yes Yes
27 2022-10-14T05:47:52.600 OTTB 282.28 24.40 6.14 7.10 317.33 -65.42 Yes Yes
28 2022-10-14T06:47:24.045 OTTB 287.80 16.05 8.99 6.58 265.49 -102.01 No Yes
29 2022-10-14T07:12:28.740 CPL 291.66 18.70 2.81 7.06 215.53 -13.67 Yes Yes
30 2022-10-14T07:17:36.155 OTTB 289.57 16.41 5.80 7.06 306.75 -18.63 Yes Yes
31 2022-10-14T09:56:04.850 CPL 308.30 5.13 5.09 7.84 149.19 -96.41 No Yes
32 2022-10-14T10:28:39.825 OTTB 290.95 26.03 6.72 7.36 139.10 -33.50 Yes Yes
33 2022-10-14T11:21:05.325 CPL 284.30 24.38 8.98 10.75 136.81 -44.76 No Yes
34 2022-10-14T11:27:32.730 CPL 293.39 19.48 3.52 30.99 44.95 -13.14 No Yes
35 2022-10-14T11:58:40.230 OTTB 284.20 15.58 5.05 9.75 85.89 -10.25 Yes Yes
36 2022-10-14T12:47:44.045 BB 301.50 27.49 3.25 29.16 52.81 -4.77 No Yes
37 2022-10-14T13:20:31.720 PL 291.35 24.31 7.06 8.90 244.57 -44.63 Yes Yes
38 2022-10-14T13:21:36.375 CPL 288.70 23.05 2.93 9.55 691.62 -16.21 Yes Yes
39 2022-10-14T13:38:25.400 OTTB 311.32 12.79 10.29 8.01 55.88 -67.71 Yes Yes
40 2022-10-14T17:35:01.675 SoftBand 278.66 31.91 9.25 7.67 47.79 -38.67 No Yes
41 2022-10-14T17:47:59.175 SoftBand 280.58 31.65 1.46 7.20 486.30 -130.42 No Yes
42 2022-10-14T19:21:39.020 OTTB 302.72 18.95 5.94 8.03 321.19 -10.00 No Yes
43 2022-10-14T20:48:47.425 SoftBand 286.06 37.31 11.61 6.97 254.97 -102.61 No Yes
44 2022-10-15T02:13:54.545 SoftBand 288.25 4.63 1.26 11.28 128.58 -17.94 Yes Yes
45 2022-10-16T10:43:17.080 BB 282.62 19.36 8.50 8.81 22.04 -3.94 No Yes
46 2022-10-17T04:43:27.720 CPL 276.95 9.82 7.64 7.45 16.37 -10.85 Yes No
47 2022-10-17T14:17:40.650 OTTB 291.35 21.34 10.91 10.52 47.56 -15.97 Yes No
48 2022-10-17T15:14:07.600 PL 280.53 12.30 9.83 6.61 268.73 -221.31 No Yes
49 2022-10-17T15:53:30.340 BB 279.93 16.88 14.97 7.50 13.36 -12.99 Yes No
50 2022-10-17T16:35:38.085 CPL 281.66 12.46 8.73 9.82 48.81 -26.24 Yes Yes
51 2022-10-18T13:57:25.965 SoftBand 286.87 2.99 10.27 7.23 4.82 -10.47 Yes Yes
52 2022-10-18T16:22:09.650 CPL 300.25 25.07 8.56 11.35 295.70 -5.49 Yes No
53 2022-10-22T04:44:52.700 CPL 300.97 29.09 3.36 21.45 86.99 -4.82 No Yes
54 2022-10-22T16:30:16.000 PL 301.02 40.42 4.85 9.18 109.45 -134.88 Yes No
55 2022-10-25T18:42:31.610 BB 277.10 4.59 12.13 7.22 14.00 -6.36 No No
56 2022-10-31T23:08:59.180 CPL 277.14 15.63 2.91 10.49 14.27 -4.90 Yes Yes
57 2022-11-09T09:05:47.515 CPL 291.45 14.72 3.44 8.74 298.00 -10.95 No No
58 2022-11-09T14:33:10.135 CPL 288.10 19.74 3.98 26.27 432.66 -7.65 Yes Yes
59 2022-11-09T15:29:02.435 CPL 286.99 18.24 3.06 8.52 502.87 -1.47 Yes No
60 2022-11-09T15:58:26.975 OTTB 288.62 26.59 15.04 7.32 71.15 -69.01 Yes Yes
61 2022-11-09T16:06:08.595 CPL 290.43 21.32 2.45 6.07 2537.76 -3.71 Yes Yes
62 2022-11-09T16:49:33.240 CPL 294.07 32.15 10.32 14.52 643.86 -14.48 No Yes
63 2022-11-09T16:50:28.385 OTTB 292.75 26.98 18.72 7.35 39.40 -27.64 No Yes
64 2022-11-09T16:51:37.330 PL 280.08 26.93 8.11 13.33 31.15 -5.87 No No
65 2022-11-09T16:52:01.750 OTTB 295.82 27.61 5.45 6.92 33.43 223.88 No No
66 2022-11-09T16:52:26.585 OTTB 278.83 23.98 9.29 6.68 111.64 -29.51 No No
67 2022-11-09T16:54:17.565 CPL 295.70 23.93 3.80 15.72 359.39 -13.91 No No
68 2022-11-09T16:55:20.665 BB 287.46 29.18 13.97 7.39 30.20 -10.96 No No
69 2022-11-09T16:57:30.950 BB 278.25 15.69 17.94 7.45 69.34 -13.18 No No
70 2022-11-09T16:59:56.580 CPL 290.30 20.83 3.57 19.85 100.90 -3.42 No No
71 2022-11-09T17:05:04.670 CPL 286.18 22.81 4.97 12.97 480.98 -1.80 Yes No
72 2022-11-09T17:21:41.635 OTTB 293.34 21.22 2.71 11.74 169.43 -8.33 Yes No
73 2022-11-09T17:35:49.100 OTTB 303.74 24.36 4.51 10.84 331.01 -29.99 Yes Yes
74 2022-11-09T17:46:18.390 BB 299.15 21.10 3.62 7.27 144.09 0.58 Yes Yes
75 2022-11-09T17:47:59.245 CPL 290.50 21.31 2.98 13.92 307.82 -7.27 Yes Yes
76 2022-11-09T18:25:56.375 CPL 289.59 26.79 2.63 17.78 96.94 -10.60 No Yes
77 2022-11-09T19:07:34.000 CPL 293.20 13.03 16.49 8.16 22.47 -9.50 Yes Yes
78 2022-11-09T19:08:04.250 SoftBand 276.47 9.37 2.08 8.84 97.96 -51.83 Yes Yes
79 2022-11-09T20:09:32.965 BB 289.05 27.24 6.08 10.98 69.20 -15.85 No No
80 2022-11-09T20:45:50.955 OTTB 284.97 21.56 4.86 8.45 223.92 -12.21 Yes Yes
81 2022-11-10T09:03:33.235 CPL 302.00 22.68 6.94 10.81 99.04 -11.88 Yes No
82 2022-11-10T14:01:58.135 OTTB 284.30 13.73 7.35 6.67 25.53 -5.10 Yes No
83 2022-11-10T18:10:56.205 OTTB 289.45 23.58 5.15 19.23 40.50 -5.87 No Yes
84 2022-11-10T23:43:45.995 CPL 285.79 13.09 4.22 12.58 43.11 -6.08 Yes No
85 2022-11-13T02:56:52.190 CPL 279.47 13.13 4.74 13.81 41.79 -7.96 No No
86 2022-11-14T11:09:19.010 BB 285.86 14.33 5.62 13.49 53.19 -4.31 No No
87 2022-11-15T17:14:34.865 CPL 294.84 21.22 2.91 10.28 181.28 -10.70 No Yes
88 2022-11-19T18:42:35.280 OTTB 284.95 12.68 1.96 15.21 6.65 -4.78 Yes Yes
89 2022-11-23T09:35:02.435 PL 294.73 7.75 14.95 7.37 4.97 -5.21 Yes Yes
90 2022-11-23T19:17:25.100 PL 311.40 24.56 8.94 6.53 2.82 -17.85 Yes Yes
91 2022-11-27T00:38:15.935 PL 310.60 35.64 10.30 7.98 8.41 -8.30 Yes Yes
92 2022-11-30T01:05:01.125 PL 299.69 17.49 4.36 7.01 8.99 -13.51 No Yes
93 2022-12-01T17:31:06.735 BB 276.56 38.99 4.63 7.40 9.78 -9.50 Yes Yes
94 2022-12-05T16:51:36.925 PL 306.92 13.65 7.23 6.94 1.94 -10.05 Yes No
95 2022-12-19T08:49:13.535 BB 284.86 20.70 17.26 7.52 6.97 -7.84 Yes No
96 2022-12-20T13:09:17.025 PL 310.28 18.36 8.90 7.38 6.70 -9.98 Yes Yes
97 2022-12-21T03:26:06.710 CPL 289.30 12.99 4.84 7.96 8.65 -8.95 Yes Yes
Refer to captionRefer to caption
Figure 1: The targeted search location (TL) of bursts. The red star marks the accurate position of SGR J1935+2154 (R.A. = 293.73°, Decl. = 21.90°).
Refer to captionRefer to caption
Figure 2: The angle separation between the targeted search location (TL) of burst and true position of SGR J1935+2154.
Refer to captionRefer to caption
Figure 3: The significance of burst v.s angle separation between the targeted search location (TL) and true position of SGR J1935+2154.
Figure 4: Representative examples of GECAM-detected bursts of SGR J1935+2154. For each burst, light curves are shown for high gain (HG, 6–300 keV). The background of each burst is shown by the red dotted line. The green vertical dotted line and blue dotted line are assessed by the Bayesian block (BB) algorithm.
3 Catalog Analysis
As shown above, we show light curves of 9 bursts in Fig. 4 to display the variety of these bursts. The trigger time of each burst is listed in Tables 3 and 4. SGR J1935+2154 bursts consist of a single peak (such as ID: 122, 138, and 161 of GECAM-B; ID: 3 and 42 of GECAM-C), multiple peaks (ID: 27 of GECAM-B; ID: 59 of GECAM-C), and intermediate bursts (ID: 126 of GECAM-B; ID: 23 of GECAM-C). The X-ray burst associated with the radio burst FRB 20221014 is ID: 138 of GECAM-B and ID: 42 of GECAM-C.
3.1 Burst Activity
The daily burst rate is defined as $N/P$, where $N$ is the observed burst number per day and $P$ represents the percentage of the effective observation time in a full day, excluding the time intervals when the Earth blocked SGR J1935+2154, or when GECAM passes through the South Atlantic Anomaly (SAA) area, during the instrument’s startup period.
We calculate the effective observation time [^8] GECAM-B orbits at low latitudes ($<30^{\circ}N/S$), while GECAM-C could cover nearly $90^{\circ}N/S$. We only calculate the observation time of GECAM-C at low latitudes in this paper since the background undergoes too violent changes to detect magnetar bursts when GECAM-C is across high latitudes. of GECAM-B/C individually and estimate the daily burst rate of GECAM-B/C, respectively (see Fig. 5). Because the effective detectable energy range of GECAM-B changed in the HG channel ($>$40 keV after October 2022), weaker bursts may not be detected compared to GECAM-C observations. Hence, given that the burst activity of GECAM-B is lower than that of GECAM-C ($>$15 keV), the daily burst rate estimated using GECAM-C data may be more reliable for assessing burst activity compared to using GECAM-B data after October 2022. A detailed study on the evolution of the effective detectable energy range change of GECAM-B will be reported in Paper II.
Figure 5: The burst history of SGR J1935+2154 observed by GECAM-B/C.
3.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data of GECAM-B (before October 2022) and GECAM-C (after October 2022) to search periodicity by the Lomb-Scargle Periodogram method (Lomb, 1976; Scargle, 1982; VanderPlas, 2018). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (false alarm probability: 0.001, significance level $\sim 3\sigma$). The peaks around the 50 day are likely caused by observation windows or bursts gap (Xie et al., 2022).
To uncover the burst activity of SGR J1935+2154 within a cycle, the start time of all bursts is folded into different phases at a given period $P=134.63$ day through the period folding method (CHIME/FRB Collaboration et al., 2020b; Zhang et al., 2021; Xie et al., 2022),
$\phi=\frac{T-T_{0}}{P}-\mathrm{floor}\biggl{(}\frac{T-T_{0}}{P}\biggr{)},$ (7)
where $\phi$ is the folded phase, $T$ is the burst start time, $T_{0}$ is the phase start point (MJD: 59215; UTC: 2021-01-01), and the floor is a function that returns the floor of the input number. As shown in Fig 7, the active day within a cycle (duty cycle) of SGR J1935+2154 bursts is 80%. The error of the period $P=134.63$ could be estimated to be 20 days with (Chime/Frb Collaboration et al., 2020),
$\sigma=\frac{PW_{\mathrm{active}}}{T_{\mathrm{span}}},$ (8)
where $W_{\mathrm{active}}$ is the active days in one cycle and $T_{\mathrm{span}}$ is the longest time separation between burst arrival times. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is $134.63\pm 20$ days, consistent with previous work (Xie et al., 2022, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
Figure 6: Lomb-Scargle periodogram of X-ray bursts from SGR J1935+2154 observed by GECAM. The vertical red dotted line indicate the peak of periodogram which is 134.63 day.
Figure 7: Phase-folded burst rate according to the 134.63 day period with MJD 59215 (2021-01-01) referenced as phase 0. The active window (i.e. from $\phi=0.7$ to $\phi=1.5$) is denoted with red dotted lines.
3.3 Burst Duration
As previous studies (Lin et al., 2013), the magnetar burst duration $T_{\mathrm{bb}}$ is calculated with the Bayesian blocks algorithm (Scargle et al., 2013). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s (at least twice the spin period of SGR J1935+2154) are treated as background. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance larger than $5\sigma$ is considered to be part of the burst region. The first block with a significance higher than $5\sigma$ is marked as the burst start time (see Tables 3 and 4). As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple peaks (e.g., burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the duration of the entire burst is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
The duration ($T_{\mathrm{bb}}$) of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We perform the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The mean values of burst duration for the four active episodes are $47.34_{-34.35}^{+125.19}$ ms for the first episode, $70.21_{-45.05}^{+125.69}$ ms for the second episode, $120.75_{-79.32}^{+231.22}$ ms for the third episode, and $85.51_{-60.77}^{+210.03}$/$83.33_{-66.32}^{+324.99}$ ms (given by GECAM-B/C data) for the fourth active episodes. The full sample gives the mean value $83.48_{-61.57}^{+234.64}$ ms. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin et al., 2020b, a; Cai et al., 2022a; Rehan & Ibrahim, 2023, 2024).
[…]
=-1cm
Table 5: The mean of burst duration distribution in each episode obtained by fitting to a lognormal function.
Instrument Episode Mean Statisticaaone-sample Kolmogorov-Smirnov test statistic p-value
GECAM-B 1st episode $47.34_{-34.35}^{+125.19}$ 0.2 0.60
2nd episode $70.21_{-45.05}^{+125.69}$ 0.07 0.95
3rd episode $120.75_{-79.32}^{+231.22}$ 0.11 0.57
4th episode $85.51_{-60.77}^{+210.03}$ 0.13 0.48
GECAM-C 4th episode $83.33_{-66.32}^{+324.99}$ 0.07 0.69
Figure 8: The burst duration of 14 X-ray bursts observed by both GECAM-B and GECAM-C.
Figure 9: The left panel is the burst duration of each burst observed by GECAM-B/C. The right panel is the burst duration distributions of GECAM-B/C. The horizontal dotted lines show the best-fit lognormal function with a mean value listed in Table 5.
3.4 Burst Waiting Time
The waiting time between the successive bursts that fall within an uninterrupted observation time interval is defined as
$\Delta{t}=t_{\mathrm{i}}-t_{\mathrm{i-1}}$ (9)
where $t_{i}$ is the burst start time of the $i-$th burst. As shown in Fig. 10, there are 45 waiting times of GECAM-B and 21 waiting times of GECAM-C during continuous observation time intervals.
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and a mean value of $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C. Compared to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B.
Figure 10: The left panel is the waiting time (time interval between two successive bursts observed by GECAM-B/C) of the X-ray bursts v.s their occurring times, where the time of the FRB 20221014 is represented by the green dashed line. The right panel is the distribution of waiting time. The horizontal dotted lines represent the best-fit lognormal function with a mean of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C, respectively.
3.5 Burst Hardness Ratio
The net counts of the source is estimated as
$N_{\mathrm{i}}=C_{\mathrm{i}}-B_{\mathrm{i}},$ (10)
where $C_{\mathrm{i}}$ and $B_{\mathrm{i}}$ represent the total counts and background counts in the burst duration, respectively. The hardness ratio is the ratio of net counts in different energy ranges.
Since changes in the effective detectable energy range of GECAM-B over the two years, we measure the hardness ratio of all bursts within the same effective range of 50-200/40-50 keV for GECAM-B. The hardness ratio of GECAM-C is calculated in the range of 30-200/15-30 keV. The hardness ratio of the four active episodes for the GECAM-B data yields mean values of $2.72\pm 1.70$, $2.25\pm 1.20$, $1.87\pm 0.63$, and $1.48\pm 0.47$, respectively. The hardness ratio of the last episode for the GECAM-C data yields a mean value of $0.76\pm 0.33$. It presents that the evolution of the burst hardness tends to be softer through the observation time, as shown in Fig. 11. More detailed spectral properties and evolution will be reported in Paper II.
Figure 11: The evolution of hardness ratio of each burst in 50-200/40-50 keV for GECAM-B and 30-200/15-30 keV for GECAM-C.
4 Summary and Discussions
In this paper, we perform a targeted sub-threshold search GECAM daily EVT data for SGR J1935+2154 bursts during the year of 2021 and 2022. We design five spectral templates to target search and find 159 X-ray bursts from GECAM-B and 97 bursts from GECAM-C (refer to Tables 3 and 4). Among the templates, CPL, OTTB, and BB cover a higher proportion of them compared to others (refer to Table 2), providing more accuracy in location (see Fig. 2). Therefore, a simple power-law function and exponential cutoff function with softer parameter settings might be preferred for searching magnetar X-ray bursts.
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, which results in the most likely period of $134.63\pm 20$ days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into four active episodes (see Fig. 5). The period of $134.63\pm 20$ days with 80% duty cycle is well consistent with previous study (Xie et al., 2022, evaluated from July 2014 to January 2022). Unlike other periods derived in previous work ($\sim 231$ days evaluated from 2014 to 2020 or $\sim 238$ days evaluated from July 2014 to October 2021 in Grossan, 2021; Zou et al., 2021, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
Table 6: Comparison of the burst duration of SGR 1935+2154 burst episodes since its discovery in 2014, as Observed by the Fermi/GBM and GECAM-B/C.
Episode Instrument Number of bursts DurationaaThe lognormal mean value of burst duration.
Jul 2014 GBM 3
Feb-Mar 2015 GBM 24 $78.00_{-14.00}^{+17.00}$
May-Jun 2016 GBM 42 $72.00_{-6.00}^{+7.00}$
Jun-Jul 2016 GBM 54 $128.00_{-10.00}^{+11.00}$
Nov 2019 GBM 22 $121.00_{-33.00}^{+45.00}$
Apr-May 2020 GBM 151 $182.00_{-19.00}^{+22.00}$
Jan 2021 GECAM-B 14 $47.34_{-34.35}^{+125.19}$
Sep 2021 GBM 79 $77.34\pm 1.03$
GECAM-B 57 $70.21_{-45.05}^{+125.69}$
Jan 2022 GBM 112 $97.03\pm 1.03$
GECAM-B 46 $120.75_{-79.32}^{+231.22}$
Oct-Dec 2022 GECAM-B 42 $85.51_{-60.77}^{+210.03}$
GECAM-C 97 $83.33_{-66.32}^{+324.99}$
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy range when both satellites detect the same burst. The burst duration of SGR J1935+2154 is around 100 ms, see Table 6 for comparison among different burst episodes. The waiting time also conforms to a lognormal function with mean values of $338.84_{-212.95}^{+573.16}$ seconds and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-B/C, respectively (refer to Fig. 10). Because of the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). The hardness ratio of X-ray bursts tends to become softer over the course of the observation time spanning these two years (see Fig. 11).
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity, as observed by Fermi/GBM (e.g., Rehan & Ibrahim, 2024). The magnetar bursts also exhibit a softer spectrum over the two-year period, as presented by the temporal evolution of the hardness ratio and a detailed spectral analysis will follow. Similar to FRB 200428, which occurred during the most active episode (April-May 2020) preceding 2020, the radio burst FRB 20221014 are detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017). Leading to such burst activities may correspond to some sudden changes in the magnetic configuration of the magnetars (e.g., starquakes, magnetic field reconnections Thompson & Duncan, 1995; Thompson & Blaes, 1998; Jones, 2003; Levin & Lyutikov, 2012). In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang & Zhang, 2021; Xie et al., 2024).
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0550300, XDA30050000), the National Key R&D Program of China (2021YFA0718500), the National SKA program of China (2020SKA0120300), the National Natural Science Foundation of China (Grant No. 12393811, 12173038, 12303045, 12273042), the Strategic Priority Research Program on Space Science of the Chinese Academy of Sciences (Grant No. XDA15360102, XDA15360300, XDA15052700), the Natural Science Foundation of Hebei Province (No. A2023205020), and the Science Research Program of Dezhou University (2024xjrc142).